### v2

# Nordhaus–Gaddum-type theorems for maximum average degree

Abstract
A $k$-decomposition $(G_{1},\dots,G_{k})$ of a graph $G$ is a partition of its edge set into $k$ spanning subgraphs $G_{1},\dots,G_{k}$. The classical theorem of Nordhaus and Gaddum bounds $\chi(G_{1})+\chi(G_{2})$ and $\chi(G_{1})\chi(G_{2})$ over all 2-decompositions of $K_{n}$.
For a graph parameter $\mathfrak{p}$}, let ${\mathfrak{p}}(k,G)=\max\{\sum_{i=1}^{k}{\mathfrak{p}}(G_{i})\}$, taken over all $k$-decompositions of graph $G$. In this paper we consider $M(k,K_{n})=M(k,n)=\max\{\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\}$, taken over all $k$-decompositions of the complete graph $K_{n}$, where $\mathrm{Mad}(G)$ denotes the maximum average degree of $G$, $\mathrm{Mad}(G)=\max\{2e(H)/|H|:H\subseteq G\}=\max\{d(H):H\subseteq G\}$.
Among the many results obtained in this paper we mention the following selected ones.
(1) $M(k,n)<\sqrt{k}\,n$, and $\lim_{k\to\infty}\bigl(\liminf_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}\bigr)=1$.
(2) Exact determination of $M(2,n)$.
(3) Exact determination of $M(k,n)$ when $k=\binom{n}{2}-t$, $0\leq t\leq(n-1)^{2}\!/3$.
A main tool used along this paper is the following list variant of Nordhaus–Gaddum, which is of independent interest and can be adopted to attack other Nordhaus–Gaddum type problems, too. A list $\mathsf{L}$ of $k$ graphs is a multiset, $\mathsf{L}=\{G_{1},\dots,G_{k}\}$, where repetitions are allowed. For a positive integer $N$ define $M^{\mathsf{L}}(k,N)=\max\{\sum\mathrm{Mad}(G_{j})\}$ taken over all lists $\mathsf{L}$ of $k$ graphs such that $\sum e(G_{j})=N$. Clearly $M(k,n)\leq M^{\mathsf{L}}(k,\binom{n}{2})$. We determine $M^{\mathsf{L}}(k,N)$ for all values of the parameters involved.
(4) Let $N\geq k\geq 2$ be any integers, and let $p,q,r$ be non-negative integers such that $N=k\binom{p}{2}+qp+r$, where $0\leq q<k$ and $0\leq r<p$. Then $M^{\mathsf{L}}(k,N)=kp-k+q$ if $r\leq(p-1)/2$, and $M^{\mathsf{L}}(k,N)=kp-k+q+1-2(p-r)/(p+1)$ if $r\geq(p-1)/2$.
Using constructions based on the theory of Steiner systems and block designs, we supply a plethora of cases where this list upper bound is indeed realized by decompositions of $K_{n}$ and equality $M(k,n)$ = $M^{\mathsf{L}}(k,\binom{n}{2})$ holds.
Applications of these bounds to other parameters considered before in the literature are given.
Keywords: Vertex degree; Maximum average degree;          Nordhaus–Gaddum theorem
AMS Subject Classification 2020: 05B05, 05C07, 05C35
1 Introduction
Nordhaus and Gaddum [20] proved that if $G$ is a graph of order $n$ then $2\sqrt{n}\leq\chi(G)+\chi(\overline{G})\leq n+1$ and $n\leq\chi(G)\cdot\chi(\overline{G})\leq(n+1)^{2}\!/4$. This theorem was the starting point of a reach theory of extremal results of the following form, called Nordhaus–Gaddum-type theorems.
A $k$-decomposition of a graph $G$ is a partition of its edge set with $k$ spanning subgraphs $G_{1},\dots,G_{k}$. For a graph parameter $\mathfrak{p}$}, this track of research investigates:
$(a)$ $\max\,\{\sum_{i=1}^{k}{\mathfrak{p}}(G_{i})\}$,
$(b)$ $\min\,\{\sum_{i=1}^{k}{\mathfrak{p}}(G_{i})\}$,
$(c)$ $\max\,\{\prod_{i=1}^{k}{\mathfrak{p}}(G_{i})\}$,
$(d)$ $\min\,\{\prod_{i=1}^{k}{\mathfrak{p}}(G_{i})\}$,
where the maxima and minima are taken over all $k$-decompositions of $G$. Analogous functions have been studied in other combinatorial structures, too.
Nordhaus and Gaddum considered the case $k=2$ for the parameter ${\mathfrak{p}}(G)=\chi(G)$, the chromatic number. Many other graphical parameters were studied in this context since then (more than 2000 articles in Google Scholar). We refer the reader to the survey [2].
Also, there is an evolving literature where Nordhaus-Gaddum type theorems are proved for $k\geq 3$; see, for example, [1], [4][6], [13], [18], [21].
In this paper we concentrate on a problem of type $(a)$, and consider the extremal function
$M(k,n):=\max\,\{\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\}$
taken over all $k$-decompositions of the complete graph $K_{n}$, where $\mathrm{Mad}(G)=\max\{2e(H)/|H|:H\subseteq G\}=\max\{d(H):H\subseteq G\}$, called the maximum average degree of $G$.
Somewhat surprisingly, we were unable to find any paper concerning Nordhaus–Gaddum-type results for the parameter $\mathrm{Mad}(G)$, despite that it is used frequently in determining the arboricity of graphs and also in many coloring problems. See, e.g., [7], [19].
The value $\mathrm{Mad}(G)+1$ dominates or is closely related to many other parameters, including the clique number $\omega(G)$, the choice number or list chromatic number $\mathrm{ch}(G)$, and the Szekeres–Wilf number—also called coloring number—denoted $\mathrm{col}(G)$, which is just one greater than the degeneracy of $G$. Hence $M(k,n)$ serves as an upper bound for the corresponding Nordhaus–Gaddum-type theorems for these parameters. We shall use this fact to comment on earlier results obtained in [13].
To deal with $M(k,n)$, we also introduce a list variant of Nordhaus–Gaddum as follows. A list $\mathsf{L}$ of k graphs is a multiset, $\mathsf{L}=\{G_{1},\dots,G_{k}\}$, where repetitions are allowed. For a positive integer $N$ define: $M^{\mathsf{L}}(k,N)=\max\{\sum\mathrm{Mad}(G_{j})\}$ taken over all lists $\mathsf{L}$ of $k$ graphs such that $\sum e(G_{j})=N$.
Clearly $M(k,n)\leq M^{\mathsf{L}}(k,\binom{n}{2})$. Hence if there is a list of $k$ graphs realizing $M^{\mathsf{L}}(k,\binom{n}{2})$ and this list can also be realized as a decomposition of $K_{n}$, then the the inequality $M(k,n)\leq M^{\mathsf{L}}(k,\binom{n}{2})$ turns to equality. This idea can be adopted for other Nordhaus–Gaddum-type results with distinct parameters.
Organization of the paper
In Section 2 we present simple properties of $\mathrm{Mad}(G)$, which are used later along the paper. In Section 3, among several results, we prove the general upper bound: $M(k,n)<\sqrt{k}\,n$ (Lemma 14 and Corollary 17). In Section 4 we deal with cases of small $k$; most importantly, we determine $M(2,n)$ exactly for all $n$ (Theorem 18). The lower-bound constructions for $M(k,n)$ for $3\leq k\leq 7$, collected in Proposition 19, are closely related to those due to Bickle [5].
In Section 5.1 we provide a complete determination of $M^{\mathsf{L}}(k,N)$ (Theorem 20) and a sufficient condition for $M(k,n)=M^{\mathsf{L}}(k,\binom{n}{2})$ (Theorem 21), which are among the major results of this paper. In the appendix we supply also a second, more algorithmic and direct proof of Theorem 21. As an application of these results, in Section 5.2 we derive the exact formula for $M(\binom{n}{2}-t,n)$ in the entire range of $0\leq t\leq(n-1)^{2}\!/3$, and characterize the extremal decompositions.
In Section 6 we present various constructions based on Steiner systems, finite geometries, Wilson’s decomposition theorem and pairwise balanced $K$-designs where $K$ is of the form $\{p,p+1\}$, to obtain sharp or nearly sharp results on various values of $M(k,n)$. Our methods can be applied to derive the inequality $(1-o_{k}(1))\sqrt{k}\,n-c_{k}<M(k,n)<\sqrt{k}\,n$ for every fixed $k$ and suitable constant $c_{k}$ (Theorem 34), concerning the asymptotic growth of $M(k,n)$. This is another major result of this paper. As $k$ grows, it means that the convergence $\lim_{k\to\infty}\bigl(\liminf_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}\bigr)=1$ is valid.
Lastly in Section 7 we consider applications of our results to several parameters dominated by or related to $M(k,n)$, which were already discussed in [13], and also several further ones. Our main result here, Theorem 40, is that if $M(k,n)=M^{\mathsf{L}}(k,\binom{n}{2})$, then the analogously defined extremal function ${\mathfrak{p}}(k,n)$ for many further graph parameters $\mathfrak{p}$} is equal to $M(k,n)$ or $\left\lfloor M(k,n)\right\rfloor+k$. This includes clique number, chromatic number, vertex/edge connectivity, and more.
Section 8 contains open problems for future research.
1.1 Notation
We apply the following standard notation for graphs $G$ :  $|G|=|V(G)|$ is the number of vertices (the order of $G$), $e(G)$ is the number of edges, $\delta(G)$ is the minimum degree of vertices, and $d(G)=(\sum_{v\in V(G)}\deg(v))/|G|={2e(G)}/{|G|}$ is the average degree.
Further common notation is $\overline{G}$ for the complementary graph of $G$, $G[X]$ for the subgraph induced by a vertex subset $X\subseteq V(G)$, and $\deg(v)$ for the degree of a vertex $v\in V(G)$.
The “support” of $G$, denoted by $G{}^{\mathsf{+}}$, is defined as the subgraph obtained from $G$ by omitting the isolated vertices. We write $n{}^{\mathsf{+}}=|G{}^{\mathsf{+}}|$ for the number of non-isolated verices, and define the “essential average degree” as $\mathrm{ed}(G)=(\sum_{v\in V(G{}^{\mathsf{+}})}\deg(v))/n{}^{\mathsf{+}}=2e(G)/n{}^{\mathsf{+}}=d(G{}^{\mathsf{+}})$, that is the average degree in $G{}^{\mathsf{+}}$.
Well known particular types of graphs are the complete graph $K_{n}$ of order $n$, and the path $P_{s}$ on $s$ vertices (having s1s-1 edges). For integers $t_{1},t_{2}>0$, the graph $t_{1}G_{1}\cup t_{2}G_{2}$ is the vertex-disjoint union of $t_{1}$ copies of $G_{1}$ and $t_{2}$ copies of $G_{2}$. We also consider multisets of graphs. The collection of $t_{1}$ copies of $G_{1}$ and $t_{2}$ copies of $G_{2}$ will be denoted by $\{t_{1}*G_{1},t_{2}*G_{2}\}$; and more generally, for multisets built upon $s$ graphs $G_{1},\dots,G_{s}$, we write $\{t_{1}*G_{1},\dots,t_{s}*G_{s}\}$.
1.2 Functions on spanning vs. non-spanning subgraphs
Recall from above that the support $G{}^{\mathsf{+}}$ of a graph $G$ is defined as the subgraph induced by the vertices of positive degrees in $G$. If $G$ is the empty graph, we define $G{}^{\mathsf{+}}=K_{1}$.
For instance, if $G=K_{4}\cup 4K_{1}$, then $G{}^{\mathsf{+}}=K_{4}$, $d(G)=12/8=1.5$, and $d(G{}^{\mathsf{+}})=3=\mathrm{ed}(G)$. As another parameter, considering edge connectivity $\lambda(G)$, for the same $G=K_{4}\cup 4K_{1}$ we have $\lambda(G)=0$ while $\lambda(G{}^{\mathsf{+}})=\lambda(K_{4})=3$.
Let $H$ be any given host graph, and $\mathfrak{p}$} any given graph invariant. Let us introduce the following functions.
(1) ${\mathfrak{p}}(k,H)=\max\{\sum{\mathfrak{p}}(G_{j})\}$ taken over all decompositions of $E(H)$ into $k$ spanning subgraphs of $H$. Note that isolated vertices are allowed in any of the $G_{j}$. If $H=K_{n}$, we write ${\mathfrak{p}}(k,n)$.
(2) ${\mathfrak{p}}{}^{\mathsf{+}}(k,H)=\max\{\sum{\mathfrak{p}}(G_{j}{}^{\mathsf{+}})\}$ taken over all decompositions of $E(H)$ into $k$ spanning subgraphs of $H$. Note that $G{}^{\mathsf{+}}$ eliminates the isolated vertices from $G$. If $H=K_{n}$, we write ${\mathfrak{p}}{}^{\mathsf{+}}(k,n)$.
A list $\mathsf{L}$ of k graphs is a multiset, $\mathsf{L}=\{G_{1},\dots,G_{k}\}$, where repetitions are allowed.
(3) For a positive integer $N$ define: ${\mathfrak{p}}^{\mathsf{L}}(k,N)=\max\{\sum{\mathfrak{p}}(G_{j})\}$ taken over all111 Interesting problems arise also where the universe of graphs is restricted, or just the admissible lists themselves are required to satisfy certain constraints. However, in this paper we deal with general lists admitting all graphs. lists $\mathsf{L}$ of $k$ graphs such that $\sum e(G_{j})=N$. It means that $N$ edges with non-specified endpoints are given, and we are asked to build $k$ graphs $G_{1},\dots,G_{k}$ from them to maximize ${\mathfrak{p}}(G_{1})+\ldots+{\mathfrak{p}}(G_{k})$.
It depends on the invariant $\mathfrak{p}$} whether or not it is sensitive with respect to the presence of isolated vertices. In the current examples $d(G)$ and $\lambda(G)$ are, but $\mathrm{Mad}(G)$ and $\mathrm{ed}(G)$ aren’t. For insensitive parameters it is not necessary to assume that all the graphs $G_{1},\dots,G_{k}$ forming a decomposition of $K_{n}$ are spanning subgraphs, the definitions can be relaxed to subgraphs. Further, if $\mathfrak{p}$} is monotone increasing in the sense that ${\mathfrak{p}}(G)\leq{\mathfrak{p}}(H)$ whenever $G\subset H$, then edge-disjoint packings may be studied instead of decompositions.
2 Some simple properties of Mad(G)
We introduce the following term, which will simplify some arguments.
Definition 1.
(Free edge.) In a graph $G$, we call an $e\in E(G)$ a free edge if its removal does not modify the $\mathrm{Mad}$ value; i.e., $\mathrm{Mad}(G-e)=\mathrm{Mad}(G)$.
Proposition 2.
(Monotonicity.) For every $n$ and $k$,
  • $(i)$
    $M(k,n+1)\geq M(k,n)+1$ if $k\leq\binom{n}{2}$.
  • $(ii)$
    $M(k+1,n)\geq M(k,n)+1/3$ if $k<\binom{n}{2}$.
Proof.
Let $G_{1},\dots,G_{k}$ decompose $K_{n}$ with $M(k,n)=\sum_{i=1}^{k}\mathrm{Mad}(G_{i})$, and let $V(K_{n+1})=V(K_{n})\cup\{w\}$. Assume that $\mathrm{Mad}(G_{k})$ is attained on a vertex set $X$, i.e., $\mathrm{Mad}(G_{k})=2e(G_{k}[X])/|X|$. Extend now $G_{k}$ with vertex $w$ and all edges $vw$ where $v\in V(K_{n})$, hence the join $G_{k}+K_{1}$. In the extended $G_{k}$ the number of edges induced by $X\cup\{w\}$ is $e(G_{k}[X])+|X|$, therefore the maximum average degree is increased to
$\frac{2e(G_{k}[X])+2|X|}{|X|+1}=\frac{2e(G_{k}[X])+(|X|-1)}{|X|+1}+1\geq\frac{2e(G_{k}[X])}{|X|}+1$
because $X$ cannot induce more than $\binom{|X|}{2}$ edges. This proves $(i)$.
For $(ii)$, let $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})=M(k,n)$, and suppose that $e(G_{1})$ is largest among the $e(G_{i})$. Then $e(G_{1})>1$, hence the number $n_{1}$ of non-isolated vertices in $G_{1}$ is at least 3. If $G_{1}$ has a free edge $e$, then taking $G_{k+1}=K_{2}$ whose edge is $e$ we increase the sum of the $\mathrm{Mad}(G_{i})$ by 1. If this is not the case, then $\mathrm{Mad}(G_{1})=2e(G_{1})/n_{1}$, each edge contributing to $\mathrm{Mad}(G_{1})$ with $2/n_{1}\leq 2/3$. Thus, picking an arbitrary edge $e$ and creating also here $G_{k+1}=K_{2}$ whose edge is $e$, $\mathrm{Mad}(G_{1})-\mathrm{Mad}(G_{1}-e)\leq 2/3$ holds, while $\mathrm{Mad}(G_{k+1})=1$. So, this modification increases the sum of the $\mathrm{Mad}$ values by at least 1/3. ∎
Comparison of the two parts of Theorem 23 provides us with $(n-1)^{2}\!/6$ values of $k$ for which the increase of 1/3 given in $(ii)$ above is tight.
Proposition 3.
If $G$ is not empty, then there is an induced connected subgraph $H\subseteq G$ realizing $\mathrm{Mad}(G)=d(H)$.
Proof.
Assume that $H\subseteq G$ satisfies $d(H)=\mathrm{Mad}(G)$, and under this condition $|V(H)|$ is smallest. It is clear that all edges induced by $V(H)$ belong to $E(H)$. Suppose for a contradiction that $H$ is disconnected, say $H=H^{\prime}\cup H^{\prime\prime}$ and there is no edge between $H^{\prime}$ and $H^{\prime\prime}$. We have $d(H^{\prime})=2e(H^{\prime})/|H^{\prime}|$, $d(H^{\prime\prime})=2e(H^{\prime\prime})/|H^{\prime\prime}|$, and $d(H)=2(e(H^{\prime})+e(H^{\prime\prime}))/(|H^{\prime}|+|H^{\prime\prime}|)$. Assuming $d(H^{\prime})\leq d(H^{\prime\prime})$, and taking their weighted average we clearly have
$d(H^{\prime})\leq\frac{|H^{\prime}|}{|H^{\prime}|+|H^{\prime\prime}|}\,d(H^{\prime})+\frac{|H^{\prime\prime}|}{|H^{\prime}|+|H^{\prime\prime}|}\,d(H^{\prime\prime})\leq d(H^{\prime\prime})\,.$
Observe that the expression in the middle is exactly $\mathrm{Mad}(H)$. Thus, $H^{\prime\prime}$ has fewer vertices than $H$ has, while its average degree is not smaller. This contradicts the choice of $H$. ∎
Proposition 4.
If $T$ is a tree, then $\mathrm{Mad}(T)=\mathrm{ed}(T)=d(T)=2-2/|T|$. More generally, if $G$ is a maximal $\delta$-degenerate graph on more than $\delta$ vertices, then $\mathrm{Mad}(G)=\mathrm{ed}(G)=d(G)=2\delta-\frac{\delta(\delta+1)}{|G|}$.
Proof.
Assume that the subgraph $G^{\prime}\subseteq G$ attains $d(G^{\prime})=\mathrm{Mad}(G)$, and $G^{\prime}$ contains no free edges. We apply the facts that a maximal $\delta$-degenerate graph on $n>\delta$ vertices has exactly $\delta n-\binom{\delta+1}{2}$ edges, moreover every subgraph of a $\delta$-degenerate graph is also $\delta$-degenerate. In this way we obtain
$\mathrm{Mad}(G)$ $=$ $d(G^{\prime})\ =\ \frac{2e(G^{\prime})}{|G^{\prime}|}\ \leq\ 2\delta-\frac{\delta(\delta+1)}{|G^{\prime}|}$
$\leq$ $2\delta-\frac{\delta(\delta+1)}{|G|}\ =\ d(G)\ =\ \mathrm{ed}(G)$
$\leq$ $\mathrm{Mad}(G)\,.$
Thus, equality holds throughout. It also follows that $G^{\prime}=G$. ∎
Concerning lower-bound constructions for $M(k,n)$ with fixed $k$, it is useful to state the following simple implication explicitly.
Proposition 5.
If $p>q>1$ and $G$ is the complete split graph $K_{p}-K_{q}$, then $\mathrm{Mad}(G)=\mathrm{ed}(G)=d(G)$.
Proof.
The graph $K_{p}-K_{q}$ is (pq)(p-q)-degenerate, and it is maximal with this property because inserting a new edge creates a subgraph $K_{p-q+2}$. Thus, Proposition 4 can be applied. ∎
Observation 6.
Since $\mathrm{Mad}(G)\geq\mathrm{ed}(G)\geq d(G)=2e(G)/|V(G)|$ holds for every graph $G$, the following inequalities are immediately seen, where the maximum is taken over all edge-disjoint packings $\mathcal{G}=\{G_{1},\dots,G_{k}\}$ of $k$ subgraphs of $K_{n}$ :
$M(k,n)=\max_{\mathcal{G}}\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\geq\max_{\mathcal{G}}\sum_{i=1}^{k}\mathrm{ed}(G_{i})\geq\max_{\mathcal{G}}\sum_{i=1}^{k}d(G_{i})\,.$
Observation 7.
To have $\mathrm{Mad}(G)=d(G)$ for a graph $G$, also $\mathrm{Mad}(G)=\mathrm{ed}(G)$ must hold. This forces $\delta(G)\geq d(G)/2$, for otherwise the removal of a vertex $v$ of smallest degree $d(v)=\delta(G)<|V(G)|/2$ would yield $d(G-v)=2(e(G)-\delta(G))/(|V(G)|-1)>(2e(G)-d(G))/(|V(G)|-1)=2e(G)/|V(G)|=d(G)$. Dense examples of graphs $G$ with $\mathrm{Mad}(G)=d(G)$ are obtained, e.g., by removing at most $n/2$ edges from $K_{n}$. Sparser examples are all regular graphs, and many more.
3 General estimates and constructions
In the first part of this section we determine the largest possible value of $\mathrm{Mad}(G)$ among graphs with a given number $m\geq 1$ of edges; that is,
$g(m):=\max\{\mathrm{Mad}(G)\mid e(G)=m\}\,.$
As one may observe, with the notation defined in the introduction, we have $g(m)=M^{\mathsf{L}}(1,m)$.
In the second part of the section we describe a structure class of edge decompositions, on which it would suffice to determine $\max\{\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\}$ in order to determine $M(k,n)$.
3.1 Single graph optimizing Mad(G)
Theorem 8.
Let $p\geq 2$ be an integer.
  • $(i)$
    If $m=\binom{p}{2}$, then g(m)=p1g(m)=p-1, with the unique extremal graph $K_{p}$.
  • $(ii)$
    If $\binom{p}{2}<m<\binom{p+1}{2}$, say $m=\binom{p}{2}+r$ where $0<r<p$, then $g(m)=\max\{p-1,2m/(p+1)\}$, attained by precisely the following graphs $G$ having $m$ edges:
    1. (a)
      any $G$ with $K_{p}\subset G$, if $0<r<\frac{p-1}{2}$;
    2. (b)
      any $G$ with $|V(G)|=p+1$, if $\frac{p-1}{2}<r<p$;
    3. (c)
      both types (a) and (b), if $r=\frac{p-1}{2}$.
Proof.
Let $G$ be any graph with $m$ edges, and $X\subseteq V(G)$ a set of vertices inducing a subgraph $H$ with $d(H)=\mathrm{Mad}(G)$. If $|X|\leq p$, then $X$ can induce at most $\binom{|X|}{2}$ edges, thus $d(H)=2e(H)/|X|\leq|X|-1\leq p-1$. If $|X|\geq p+1$, then $d(H)=2e(H)/|X|\leq 2e(G)/(p+1)$.
Both bounds are attained by graphs of $m$ edges in the given range, namely by $K_{p}$ supplemented with arbitrarily positioned $m-\binom{p}{2}$ further edges, and by all graphs with $p+1$ vertices and $m$ edges, respectively. ∎
Definition 9.
(Family of extremal (p,r)-graphs.) For $p\geq 2$ and $p\geq r\geq 0$, let $m=\binom{p}{2}+r$. Based on the above, we denote:
[…]
Definition 10.
(Representative.) For $r=0$ the representative $G_{p,\,0}$ of $\mathcal{F}_{p,\,0}$ is $K_{p}$. For any $p\geq r\geq 1$, we denote by $G_{p,\,r}$ the graph with $p+1$ vertices, obtained from $K_{p}$ by supplementing it with a further vertex of degree $r$. We view this $G_{p,\,r}$ as the representative of the family $\mathcal{F}_{p,\,r}$ of extremal graphs for $g(m)$ where $m=\binom{p}{2}+r$. In particular, $G_{p,\,p}\cong K_{p+1}$.
As in the definition of representative we allow both $r=0$ and $r=p$, we can view $K_{p+1}$ as $G_{p+1,\,0}$ a representative of $\mathcal{F}_{p+1,\,0}$ and also as $G_{p,\,p}$ a representative of $\mathcal{F}_{p,\,p}$, where the two families coincide.
One may also note that, for 0<r<(p1)/20<r<(p-1)/2, the $r$ edges incident with the low-degree vertex of $G_{p,\,r}$ are free edges (and the others are not).
3.2 Equivalent condition for M (k , n)
Concerning the next result, let us recall that the graphs $G_{1},\dots,G_{k}$ attaining $M(k,n)=\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$ do not necessarily decompose $K_{n}$, some $e\in E(K_{n})$ may not belong to any of those $G_{i}$.
Proposition 11.
Let $n\geq 3$ and $k\geq 2$. Suppose that the packing of graphs $G_{1},\dots,G_{k}$ optimzes $M(k,n)$. Let $X_{j}\subset V(G_{j})$, $j=1,\dots,k$ be a subset on which $\mathrm{Mad}(G_{j})$ is realized, namely $2e(G_{j}[X_{j}])/|X_{j}|=\mathrm{Mad}(G_{j})$.
Assume $|X_{1}|\leq\dots\leq|X_{k}|$. Then, taking $H_{1}$ as the complete graph on $X_{1}$ and $V(H_{j})=X_{j}$, $E(H_{j})=E(K_{n}[X_{j}])\diagdown(\,\cup_{i<j}E(H_{i}))$ for $j\geq 2$, also the graphs $H_{1},\dots,H_{k}$ optimize $M(k,n)$ such that $\mathrm{Mad}(H_{j})$ is realized on $X_{j}$.
Proof.
To simplify notation, let us write $\mathcal{E}=E(G_{1})\cup\cdots\cup E(G_{k})$ and $\mathcal{X}=\binom{X_{1}}{2}\cup\cdots\cup\binom{X_{k}}{2}$, both viewed as subsets of $E(K_{n})$.
If $e\in E(K_{n})\diagdown\mathcal{X}$, then $e$ plays no role in $M(k,n)$, because $\mathrm{Mad}(G_{j})=\mathrm{Mad}(G_{j}[X_{j}])$ for every $j$. Hence, we may assume $\mathcal{E}\subseteq\mathcal{X}$ without loss of generality. On the other hand, if $e\in\mathcal{X}\diagdown\mathcal{E}$ held for an edge $e$, then there would exist $j$ with $e\in\binom{X_{j}}{2}$ and re-defining $E(G_{j}):=E(G_{j})\cup\{e\}$ would increase $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})$ by $2/|X_{j}|$, contradicting the assumption that $G_{1},\dots,G_{k}$ realizes $M(k,n)$. Thus $\mathcal{X}\subseteq\mathcal{E}$, consequently $\mathcal{X}=\mathcal{E}$.
We will prove that the re-distribution of the edges belonging to $\mathcal{X}$ by the rules $E(H_{1})=\binom{X_{1}}{2}$ and $E(H_{j})=\binom{X_{j}}{2}\diagdown(\,\cup_{i<j}\binom{X_{i}}{2})$ for $j\geq 2$ yields $\mathrm{Mad}(H_{1})+\ldots+\mathrm{Mad}(H_{k})\geq\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$, which clearly implies that also $H_{1},\dots,H_{k}$ optimize $M(k,n)$. For this purpose the following facts can be observed for any $1\leq i<j\leq k$ and any edge $e\in\binom{X_{i}}{2}\cap\binom{X_{j}}{2}$.
  • If $|X_{i}|<|X_{j}|$, then $e\in E(G_{j})$ is impossible, because in that case moving $e$ from $G_{j}$ to $G_{i}$ would increase $\mathrm{Mad}(G_{i})+\mathrm{Mad}(G_{j})$ with $2/|X_{i}|-2/|X_{j}|>0$, a contradiction to $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})=M(k,n)$.
  • If $|X_{i}|=|X_{j}|$, then for $e$ it is irrelevant whether $e\in E(G_{i})$ or $e\in E(G_{j})$, because in either case the contribution of $e$ to $\mathrm{Mad}(G_{i})+\mathrm{Mad}(G_{j})$ is $2/|X_{i}|=2/|X_{j}|$. (If moving $e$ from $G_{j}$ to $G_{i}$ modifies $G_{j}$ to a graph in which $X_{j}$ does not represent $\mathrm{Mad}$ anymore, it means that $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$ is improvable, contradicting that $G_{1},\dots,G_{k}$ optimize $M(k,n)$.)
The first property implies that, already at the beginning, each $e\in\mathcal{X}$ belongs to a $G_{j}$ such that $|X_{j}|=\min\{|X_{i}|\mid i\in\mathcal{I}_{e}\}$. And then, if this minimum is attained for more than one subscript $j$, the second property yields that we are free to choose any such $j$, specifically we can prescribe $j=\min\{i\in\mathcal{I}_{e}\}$ for $e$ because $|X_{1}|\leq\cdots\leq|X_{k}|$. Applying this “shift to the left” (or First Fit) rule, $j=\min\{i\in\mathcal{I}_{e}\}$ for all $e\in\mathcal{X}$ exactly means for the obtained graphs $H_{1},\dots,H_{k}$ that $H_{1}$ is the complete graph on the vertex set $X_{1}$, and for every $j>1$ the graph $H_{j}$ contains exactly those edges from $\binom{X_{j}}{2}$ which do not occur in $H_{1}\cup\cdots\cup H_{j-1}$. ∎
It is also valid that the $\mathrm{Mad}$-realizing sets for an optimal decomposition cover the entire $V(K_{n})$.
Proposition 12.
Let $n\geq 3$ and $k\geq 2$. Suppose that the packing of graphs $G_{1},\dots,G_{k}$ optimzes $M(k,n)$, and the sets $X_{j}\subset V(G_{j})$ for $j=1,\dots,k$ realize $\mathrm{Mad}(G_{j})$, i.e., $2e(G_{j}[X_{j}])/|X_{j}|=\mathrm{Mad}(G_{j})$. Then $X_{1}\cup\cdots\cup X_{k}=V(K_{n})$.
Proof.
Otherwise choose any vertex $v\in V(K_{n})\diagdown(X_{1}\cup\cdots\cup X_{k})$. The edges $vx\in E(K_{n})$ for $x\in X_{1}$ have no contribution to $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$, therefore we can replace $G_{1}$ with the graph $H:=(X_{1}\cup\{v\},E(G_{1}[X_{1}]\cup\{vx\mid x\in X_{1}\})$ to obtain a modified packing of $k$ subgraphs in $K_{n}$. Then, denoting $e_{1}=e(G_{1}[X_{1}])$ and $x_{1}=|X_{1}|$, and observing that $e_{1}\leq\frac{1}{2}(x_{1}^{2}-x_{1})$, we obtain
$\mathrm{Mad}(H)-\mathrm{Mad}(G_{1})=\frac{2e_{1}+2x_{1}}{x_{1}+1}-\frac{2e_{1}}{x_{1}}=\frac{2x_{1}^{2}-2e_{1}}{x_{1}^{2}+x_{1}}\geq 1\,.$
This contradicts the assumption that $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$ attains the maximum, $M(k,n)$. ∎
From Propositions 11 and 12 we also obtain:
Corollary 13.
The determination of $M(k,n)$ is equivalent to the following optimization problem:
$\sum_{i=1}^{k}\frac{2e(G_{i})}{|X_{i}|}\qquad\longrightarrow\qquad\max$
under the following conditions:
  • $X_{1}\cup\cdots\cup X_{k}=V(K_{n})$;
  • $V(G_{i})=X_{i}$ for $i=1,\dots,k$;
  • $E(G_{j})=\binom{X_{j}}{2}\diagdown\,{\displaystyle\cup_{i<j}}\binom{X_{i}}{2}$.
An advantage of this formulation is that it admits an integer programming approach, which may be efficient when $k$ and $n$ are not too large.
3.3 Upper bound for M (k , n) from continuous relaxation
Lemma 14.
Suppose $x_{1},\dots,x_{k}>0$ are reals such that $\sum_{j=1}^{k}x_{j}=\binom{n}{2}$, and let $y_{1},\dots,y_{k}>0$ be the reals with $y_{j}(y_{j}-1)/2=x_{j}$. Then:
  • $(1)$
    $\max\,\{\,\sum_{j=1}^{k}(y_{j}-1)\}$ is obtained when all $x_{j}$ are equal, namely $x_{j}=\binom{n}{2}/k$ for $j=1,\dots,k$.
  • $(2)$
    $\max\,\{\,\sum_{j=1}^{k}(y_{j}-1)\}=(\sqrt{k^{2}+4kn^{2}-4kn}-k)/2<\sqrt{k}\,n$.
  • $(3)$
    If $k=\binom{n}{2}/\binom{t}{2}$, then $y_{j}(y_{j}-1)/2=t(t-1)/2$ and the maximum equals k(t1)=n(n1)/tk(t-1)=n(n-1)/t.
Proof.
The definition of $y_{j}$ yields $y_{j}^{2}-y_{j}-2x_{j}=0$, the positive root is $y_{j}=\frac{1}{2}(1+\sqrt{8x_{j}+1}\,)$. Since $a\sqrt{x}+b$ is a concave function for $a>0$, the assertion of (1) follows by Jensen’s inequality.
Plugging in $x=\binom{n}{2}/k$ for all $x_{j}$, we obtain y(y1)/2=n(n1)/2ky(y-1)/2=n(n-1)/2k, hence $ky^{2}-ky-n^{2}+n=0$ and $y=(k+\sqrt{k^{2}+4kn^{2}-4kn})/2k$ for all $y_{j}$, therefore
$k(y-1)$ $=$ $k((k+\sqrt{k^{2}+4kn^{2}-4kn}-2k)/2k$
$=$ $(\sqrt{k^{2}+4kn^{2}-4kn}-k))/2$
$<$ $\sqrt{k}\,n$
since the last two lines are positive and eliminating the first square root by squaring we get $4kn^{2}+k^{2}-4kn\leq(2\sqrt{k}\,n+k)^{2}=4kn^{2}+4k^{3/2}\,n+k^{2}$, which is obviously true. This proves (2).
Finally, (3) is a consequence of (1). Indeed, by (1) the maximum of $\sum_{j=1}^{k}(y_{j}-1)$ is attained when $\binom{y_{j}}{2}=x_{j}=\binom{n}{2}/k=\binom{t}{2}$ holds for all $j=1,\dots,k$ ; consequently $y_{j}=t$ and
$(y_{1}-1)+\dots+(y_{k}-1)=(t-1)\cdot k=(t-1)\binom{n}{2}/\binom{t}{2}=n(n-1)/t\,.$
Remark 15.
Literally the same computation yields:
$M^{\mathsf{L}}(k,N)<\sqrt{2kN}\,.$
It is worth noting that the jump of estimate in the last step yielding $\sqrt{k}\,n$ in the proof of Lemma 14 is as large as $cn^{2}$ when $k$ is proportional to $n^{2}$, hence (2) cannot be within $(1+o(1))$ to the optimum if $k$ is quadratic in $n$.
In the same flavour, the following real relaxation of Theorem 8 can also be stated.
Proposition 16.
For every $m>0$, $g(m)\leq y-1$, where $y>0$ is the real number satisfying y(y1)/2=my(y-1)/2=m.
Proof.
Let $m=\binom{p}{2}+r$, where $p$ and $r$ are integers, and $0\leq r<p$. We have seen that $g(m)=\max\{p-1,2m/(p+1)\}$ holds.
If g(m)=p1g(m)=p-1, then $g(m)+1=p\leq y$ because $\binom{p}{2}\leq m=\binom{y}{2}$, with equality if and only if $r=0$.
In the other case we have g(m)=2m/(p+1)=y(y1)/(p+1)<y1g(m)=2m/(p+1)=y(y-1)/(p+1)<y-1 with strict inequality, because $\binom{y}{2}=m<\binom{p+1}{2}$ and therefore $y<p+1$. ∎
Corollary 17.
$M(k,n)\leq M^{\mathsf{L}}(k,N)<\sqrt{2kN}<\sqrt{k}\,n$, where $N=\binom{n}{2}$.
Proof.
Let $G_{1},\dots,G_{k}$ be an edge decomposition of $K_{n}$, and let $y_{j}(y_{j}-1)=2e(G_{j})$ for $j=1,\dots,k$. Applying Proposition 16 and part (2) of Lemma 14 we obtain: $M(k,n)=\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\leq\sum_{j=1}^{k}(y_{j}-1)<\sqrt{k}\,n$. The intermediate $\sqrt{2kN}$ comes from Remark 15. ∎
Later in Theorem 34 we shall see that the coefficient $\sqrt{k}$ of $n$ is very close to be best possible, for every $k$.
4 M (k , n) for small k
In this section we consider the cases $2\leq k\leq 7$. The exact value of $M(2,n)$ will be determined for all $n\geq 3$, and lower bounds will be given for the other values of $k$.
4.1 Exact formula for k = 2
Theorem 18.
If $G$ is any graph of order $n\geq 3$, then
$\mathrm{Mad}(G)+\mathrm{Mad}(\overline{G})\leq\begin{cases}\begin{tabular}[]{ll}$(5n^{2}-6n+1)/4n$&if\/ $n$ is odd,\\ $(5n^{2}-6n)/4n$&if\/ $n$ is even.\end{tabular}\end{cases}$
Moreover, both bounds are tight, hence the above formulas give the right value $M(2,n)=5n/4-3/2+\epsilon_{n}$, with $\epsilon_{n}=0$ if $n$ is even and $\epsilon_{n}=\frac{1}{4n}$ if $n$ is odd, for every $n\geq 3$.
Proof.
Let $X$ and $Y$ be subsets of $V(K_{n})$, such that $\mathrm{Mad}(G)=d(G[X])$ and $\mathrm{Mad}(\overline{G})=d(\overline{G}[Y])$. Say, $|X|=x$ and $|Y|=y$. We may assume $x\leq y$ without loss of generality, as we are free to switch between $G$ and $\overline{G}$ if necessary. Also, from Corollary 12 applied for $k=2$ (that is, $G_{1}=G$ and $G_{2}=\overline{G}$), we see that $\mathrm{Mad}(G)+\mathrm{Mad}(\overline{G})$ can be largest only if $X\cup Y=V(K_{n})$, hence $x+y\geq n$ and $|X\cap Y|=x+y-n$. Moreover, by Proposition 11, we may restrict attention to the case $E(G[X])=\binom{X}{2}$ and $E(\overline{G}[Y])=\binom{Y}{2}\diagdown\binom{X\cap Y}{2}$, whereas the edges (if any) connecting $X\diagdown Y$ with $Y\diagdown X$ can be distributed arbitrarily between $G$ and $\overline{G}$ and do not have any contribution to $\mathrm{Mad}(G)+\mathrm{Mad}(\overline{G})$. Consequently,
$\mathrm{Mad}(G)+\mathrm{Mad}(\overline{G})$ $\leq$ $x-1+\frac{y(y-1)-(x+y-n)(x+y-n-1)}{y}$
$=$ $\frac{2nx+2ny+x-x^{2}-y-xy-n^{2}-n}{y}$
$=$ $2n-1-x+\frac{2nx+x-x^{2}-n^{2}-n}{y}$
$\leq$ $2n-1-x+\frac{2nx+x-x^{2}-n^{2}-n}{n}$
$=$ $n-2+\frac{x(n+1-x)}{n}$
$\leq$ $n-2+\frac{1}{n}\left\lfloor\frac{(n+1)^{2}}{4}\right\rfloor,$
which is equivalent to the stated upper bound. Tightness is shown by the graph $G=K_{\left\lfloor(n+1)/2\right\rfloor}$ or $G=K_{\left\lceil(n+1)/2\right\rceil}$ (which coincide if $n$ is odd) and its complement. ∎
4.2 Lower-bound constructions for small k
In the following small cases of $k$ we describe two constructions, which are substantially different. Although the corresponding sums of the $\mathrm{Mad}(G_{i})$ are not exactly the same, the differences are marginal. However, there seems to be no smooth transition between the two structures presented for the same $k$. This fact may be viewed as an indication that the precise determination of $M(k,n)$ for these values of $k$ may be not at all easy.
Proposition 19.
There exist constants $c_{3},\dots,c_{7}$ such that
  • $M(3,n)\geq 3n/2-c_{3}$,
  • $M(4,n)\geq 5n/3-c_{4}$,
  • $M(5,n)\geq 11n/6-c_{5}$,
  • $M(6,n)\geq 2n-c_{6}$,
  • $M(7,n)\geq 7n/3-c_{7}$.
Proof.
Let us recall from Proposition 11 that there is a natural two-way correspondence between $\mathrm{Mad}$-optimal edge decompositions $G_{1},\dots,G_{k}$ of $K_{n}$ and sequences $X_{1},\dots,X_{k}$ of subsets of $V(K_{n})$ arranged in increasing order of cardinalities. More explicitly, the graphs $G_{j}$ can be obtained by the recursive rule $E(G_{j})=\binom{X_{j}}{2}\diagdown(\,\cup_{i<j}\,E(G_{i})\,)$.
In each case of $k=3,4,5,6$ we create two sequences $G_{1},\dots,G_{k}$. In one of them we begin with three sets $A_{1},A_{2},A_{3}$, mutually disjont, of respective cardinalities $\left\lfloor n/3\right\rfloor\leq a_{1}\leq a_{2}\leq a_{3}\leq\left\lceil n/3\right\rceil$, with $a_{1}+a_{2}+a_{3}=n$. In the other, which is usually better with a marginal extent, we begin with two disjont sets $B_{1},B_{2}$ of cardinalities $|B_{1}|=b_{1}=\left\lfloor n/2\right\rfloor$ and $|B_{2}|=b_{2}=\left\lceil n/2\right\rceil$.
$k=3$ : Set $X_{1}=A_{1}\cup A_{2}$, $X_{2}=A_{1}\cup A_{3}$, $X_{3}=A_{2}\cup A_{3}$. Then $G_{1}\cong K_{a_{1}+a_{2}}$, $G_{2}\cong K_{a_{1}+a_{3}}-K_{a_{1}}$, $G_{3}\cong K_{a_{2},a_{3}}$; hence $\mathrm{Mad}(G_{1})=a_{1}+a_{2}-1\sim 2n/3$, $\mathrm{Mad}(G_{2})\sim n/2$, $\mathrm{Mad}(G_{3})\sim n/3$, summing up to $3n/2$ approximately.
Alternatively, set $X_{1}=B_{1}$, $X_{2}=B_{2}$, $X_{3}=B_{1}\cup B_{2}=V(K_{n})$. Then $G_{1}\cong K_{\left\lfloor n/2\right\rfloor}$, $G_{2}\cong K_{\left\lceil n/2\right\rceil}$ (vertex-disjoint), and $G_{3}\cong K_{\left\lfloor n/2\right\rfloor,\left\lceil n/2\right\rceil}$. Here each $G_{i}$ has average degree near to $n/2$.
$k=4$ : Let $X_{i}=A_{i}$ for $i=1,2,3$ (hence $G_{i}\cong K_{a_{i}}$), and $X_{4}=V(K_{n})$ (hence $G_{4}\cong K_{a_{1},a_{2},a_{3}}$). Then $\mathrm{Mad}(G_{1})+\mathrm{Mad}(G_{2})+\mathrm{Mad}(G_{3})=n-3$, and $\mathrm{Mad}(G_{4})\sim 2n/3$. Interestingly, the sets $A_{i}$ also offer the solution $X_{1}=A_{1}$, $X_{2}=A_{1}\cup A_{2}$, $X_{3}=A_{1}\cup A_{3}$, $X_{4}=A_{2}\cup A_{3}$. Then $G_{1}\cong K_{a_{1}}$, $G_{2}\cong K_{a_{1}+a_{2}}-K_{a_{1}}$, $G_{3}\cong K_{a_{1}+a_{3}}-K_{a_{1}}$, $G_{4}\cong K_{a_{2},a_{3}}$. The corresponding respective average degrees are about $n/3$, $n/2$, $n/2$, $n/3$.
The other approach starts with $X_{1}=B_{1}$, $X_{2}=B_{2}$, i.e., $G_{1}\cong K_{b_{1}}$ and $G_{2}\cong K_{b_{2}}$, hence $\mathrm{Mad}(G_{1})+\mathrm{Mad}(G_{2})=n-2$. Then split $X_{2}$ into two parts, $X_{2}=Y\cup Y^{\prime}$, $|Y|=\left\lfloor b_{2}/2\right\rfloor$, $|Y^{\prime}|=\left\lceil b_{2}/2\right\rceil$; and set $X_{3}=X_{1}\cup Y$, $X_{4}=X_{1}\cup Y^{\prime}$. The obtained graphs are $G_{3}\cong K_{b_{1},\left\lfloor\left\lfloor b_{2}/2\right\rfloor/2\right\rfloor}$ and $G_{4}\cong K_{b_{1},\left\lceil\left\lfloor b_{2}/2\right\rfloor/2\right\rceil}$, each having its average degree nearly $n/3$.
$k=5$ : Let $X_{1}=A_{1}$, $X_{2}=A_{2}$, $X_{3}=A_{1}\cup A_{2}$, $X_{4}=A_{1}\cup A_{3}$, $X_{5}=A_{2}\cup A_{3}$. This yields $G_{1}\cong K_{a_{1}}$, $G_{2}\cong K_{a_{2}}$, $G_{3}\cong K_{a_{1},a_{2}}$, $G_{4}\cong K_{a_{1}+a_{3}}-K_{a_{1}}$, $G_{5}\cong K_{a_{2},a_{3}}$. The average degrees are nearly $n/3$, $n/3$, $n/3$, $n/2$, and $n/3$, respectively.
Also here we can alternatively start with $X_{1}=B_{1}$, $X_{2}=B_{2}$, hence generating $\mathrm{Mad}(G_{1})+\mathrm{Mad}(G_{2})=n-2$. For the other three graphs we cut both $B_{1}$ and $B_{2}$ into half, say $B_{1}=B_{1}^{\prime}\cup B_{1}^{\prime\prime}$ and $B_{2}=B_{2}^{\prime}\cup B_{2}^{\prime\prime}$, each of those four sets having cardinality $\left\lfloor n/4\right\rfloor$ or $\left\lceil n/4\right\rceil$. Then set $X_{3}=B_{1}^{\prime}\cup B_{2}^{\prime}$, $X_{4}=B_{1}^{\prime\prime}\cup B_{2}^{\prime}$, $X_{5}=B_{1}\cup B_{2}^{\prime\prime}$. The average degree in $G_{3}$ and in $G_{4}$ is about $n/4$, and in $G_{5}$ it is about $n/3$.
$k=6$ : Let $X_{i}=A_{i}$ for $i=1,2,3$ and $X_{4}=A_{1}\cup A_{2}$, $X_{5}=A_{1}\cup A_{3}$, $X_{6}=A_{2}\cup A_{3}$ The six graphs obtained in this way are $K_{a_{1}}$, $K_{a_{2}}$, $K_{a_{3}}$, $K_{a_{1},a_{2}}$, $K_{a_{1},a_{3}}$, $K_{a_{2},a_{3}}$. Each of them has average degree about $n/3$.
For an alternative, also here we cut both $B_{1}$ and $B_{2}$ into half: $B_{1}=B_{1}^{\prime}\cup B_{1}^{\prime\prime}$ and $B_{2}=B_{2}^{\prime}\cup B_{2}^{\prime\prime}$. Then set $X_{1}=B_{1}$, $X_{2}=B_{2}$, $X_{3}=B_{1}^{\prime}\cup B_{2}^{\prime}$, $X_{4}=B_{1}^{\prime\prime}\cup B_{2}^{\prime}$, $X_{5}=B_{1}^{\prime}\cup B_{2}^{\prime\prime}$, $X_{6}=B_{1}^{\prime\prime}\cup B_{2}^{\prime\prime}$. The average degrees are about $n/2$, $n/2$, $n/4$, $n/4$, $n/4$, $n/4$.
$k=7$ : This construction is obtained using the Fano plane, which is the first nontrivial particular case of Corollary 33 $(i)$. We will give all details there. Currently we do not know whether there is a substantially different alternative decomposition attaining $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{7})\sim 7n/3$. ∎
5 Exact formula for ML(k , N ) and sufficient condition for its equality with M (k , n)
5.1 Tight relations: Steiner systems and graph designs
We begin with a complete determination of $M^{\mathsf{L}}(k,N)$.
Theorem 20.
For any integers $N\geq k\geq 2$, let $p\geq 2$ be the unique integer for which $k\binom{p}{2}\leq N<k\binom{p+1}{2}$ holds. Write $N$ in the form $N=k\binom{p}{2}+qp+r$, where $0\leq q<k$ and $0\leq r<p$. Then
[…]
Proof.
Summing up the relevant values in the middle column of the following table (depending on $r$), the multiset $\{q*K_{p+1},\,(k-q)*K_{p}\}$ for $r=0$ or $\{q*K_{p+1},\,(k-q-1)*K_{p},\,1*G_{p,r}\}$ for $0<r<p$ shows that $M^{\mathsf{L}}(k,N)$ is at least as large as given in the theorem. We will prove that no list can have a larger $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})$ under the condition $\sum_{i=1}^{k}e(G_{i})\leq N$.
   graphs $H_{i}$ total of $\mathrm{Mad}(H_{i})$ side condition
$q$ copies of $K_{p+1}$ qpqp
kq1k-q-1 copies of $K_{p}$ (kq1)(p1)(k-q-1)(p-1)
11 copy of $G_{p,r}$ p1p-1 $r\leq(p-1)/2$
11 copy of $G_{p,r}$ p2(pr)/(p+1)p-2(p-r)/(p+1) r>(p1)/2r>(p-1)/2
Among the possibly many lists $\mathsf{L}=\{G_{1},\dots,G_{k}\}$ of $k$ graphs verifying $M^{\mathsf{L}}(k,N)$, let us choose one satisfying the following hierarchic set of extremal conditions:
  • (0)
    $e(G_{1})+\ldots+e(G_{k})\leq N$ and $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})=M^{\mathsf{L}}(k,N)$.
  • (1)
    $\mathsf{L}$ maximizes the number $s(\mathsf{L})$ of complete graphs, under (0).
  • (2)
    $\mathsf{L}$ minimizes the difference $d(\mathsf{L})$ between the maximum order and the minimum order of complete graphs in the list, under (0) and (1).
  • (3)
    $\mathsf{L}$ minimizes the number $h(\mathsf{L})$ of pairs that realize $d(\mathsf{L})$, under (0), (1), and (2).
In addition we assume the following, which can be satisfied without loss of generality as we optimize a list for $M^{\mathsf{L}}(k,N)$ :
  • (4)
    Each $G_{i}\in\mathsf{L}$ is of the form $G_{i}\cong G_{p_{i},\,r_{i}}$ with $0\leq r_{i}<p_{i}$, hence the representative of an $\mathcal{F}_{p_{i},\,r_{i}}$, having $\mathrm{Mad}(G_{i})=g(m)$ where $m=e(G_{i})=\binom{p_{i}}{2}+r_{i}$.
We are going to prove that under these conditions there exists a list $\mathsf{L}$ with $k-1\leq s(\mathsf{L})\leq k$ and $d(\mathsf{L})\leq 1$. Moreover, the possible unique non-complete member of $\mathsf{L}$ is in tight relation with the k1k-1 complete members.
There are two types of non-complete graphs $G\in\mathcal{F}_{p,\,r}$ concerning $\mathrm{Mad}(G)$, namely type A: $0<r\leq(p-1)/2$, and type B: (p1)/2<r<p(p-1)/2<r<p. However, if a $G\in\mathcal{F}_{p,\,r}$ of type A occurs in $\mathsf{L}$, then it contains $r$ free edges. Removing those edges, condition (0) is still respected, while the number $s(\mathsf{L})$ of complete members of $\mathsf{L}$ is increased, a contradiction to (1). Consequently, if there are any non-complete graphs in the list, all of them are of type B.
Suppose for a contradiction that $\mathsf{L}$ contains two members of type B, say $G_{i}=G_{p_{i},\,r_{i}}$ and $G_{j}=G_{p_{j},\,r_{j}}$, and assume $p_{i}\leq p_{j}$. We observe that $p_{i}=p_{j}$ must hold, for otherwise re-defining $G_{i}:=G_{p_{i},\,r_{i}+1}$ and $G_{j}:=G_{p_{j},\,r_{j}-1}$ would keep $e(G_{i})+e(G_{j})$ unchanged, while $\mathrm{Mad}(G_{i})+\mathrm{Mad}(G_{j})$ would increase with $\frac{2}{p_{i}+1}-\frac{2}{p_{j}+1}>0$, contradicting (0). So, let $p=p_{i}=p_{j}$, and suppose $p>r_{j}\geq r_{i}>(p-1)/2$. We now re-define $G_{j}:=K_{p+1}$ and $G_{i}:=G_{p_{i},\,r_{i}+r_{j}+1-p}$. This keeps $e(G_{i})+e(G_{j})$ unchanged, does not decrease $\mathrm{Mad}(G_{i})+\mathrm{Mad}(G_{j})$ (in fact to avoid increase and hence a contradiction to (0) we must have $r_{i}+r_{j}\geq 3(p-1)/2$), but $G_{j}$ becomes a complete graph, and so $s(\mathsf{L})$ increaes, contradicting (1).
Consequently, $k-1\leq s(\mathsf{L})\leq k$ holds. Next, we prove $d(\mathsf{L})\leq 1$. Suppose for a contradiction that $G_{i}=K_{p_{i}}$ is a smallest clique in $\mathsf{L}$, and $G_{j}=K_{p_{j}}$ is a largest clique in $\mathsf{L}$, with $p_{j}>p_{i}+1$. Let us re-define $G_{i}=K_{p_{i}+1}$ and $G_{j}=K_{p_{j}-1}$. This keeps $\mathrm{Mad}(G_{i})+\mathrm{Mad}(G_{j})$ unchanged, and decreases $e(G_{i})+e(G_{j})$ by $p_{j}-p_{i}-1>0$, so that (0) is respected. Also, the number $s(\mathsf{L})$ of complete members in $\mathsf{L}$ remains the same, respecting (1). Should $G_{i}$ be unique smallest or $G_{j}$ unique largest in $\mathsf{L}$, the modification would decrease $d(\mathsf{L})$, contradicting (2). In any other case the value of $h(\mathsf{L})$ decreases, contradicting (3).
In this way, taking the fact $N<k\cdot\binom{p+1}{2}$ into account, one of the following lists is obtained:
  1. 1.
    $\{k*K_{p^{\prime}}\}$, with $p^{\prime}\leq p$ ;
  2. 2.
    $\{q^{\prime}*K_{p^{\prime}+1},\,(k-q^{\prime})*K_{p^{\prime}}\}$, with $p^{\prime}\leq p$ and if $p^{\prime}=p$ then $q^{\prime}\leq q$ ;
  3. 3.
    $\{(k-1)*K_{p^{\prime}},\,1*G_{p_{k},\,r_{k}}\}$, with $p^{\prime}\leq p+1$ ;
  4. 4.
    $\{q^{\prime}*K_{p^{\prime}+1},\,(k-q^{\prime}-1)*K_{p^{\prime}},\,1*G_{p_{k},\,r_{k}}\}$.
The first and second cases obviously respect the formula stated in the theorem as an upper bound. The proof for the third and fourth cases are completed by inspecting the relation between $p^{\prime}$ and $p_{k}$ as it depends on whether $q^{\prime}=0$ or $q^{\prime}>0$.
We claim that $q^{\prime}=0$ implies $p_{k}=p^{\prime}$ or $p_{k}=p^{\prime}-1$. Indeed, the assumption means that the first k1k-1 graphs all are copies of $K_{p^{\prime}}$. If $p_{k}\leq p^{\prime}-2$, we move one edge from $G_{1}$ to $G_{k}$, hence re-define $G_{1}:=G_{p^{\prime}-1,\,p^{\prime}-2}$ and $G_{k}:=G_{p_{k},\,r_{k}+1}$. This keeps $e(G_{1})+e(G_{k})$ unchanged, but increases $\mathrm{Mad}(G_{1})+\mathrm{Mad}(G_{k})$ with $\frac{2}{p_{k}+1}-\frac{2}{p^{\prime}}>0$, a contradiction to (0). On the other hand, if $p_{k}>p^{\prime}$, re-define $G_{1}:=K_{p^{\prime}+1}$ and $G_{k}:=G_{p_{k}-1,\,r_{k}}$. This decreases $e(G_{1})+e(G_{k})$ with $p_{k}-1-p^{\prime}\geq 0$. Further, $\mathrm{Mad}(G_{1})$ is increased by 1, and $\mathrm{Mad}(G_{k})$ is decreased from $\frac{p_{k}(p_{k}-1)+2r_{k}}{p_{k}+1}=p_{k}-2+\frac{2r_{k}+2}{p_{k}+1}$ to $\frac{(p_{k}-1)(p_{k}-2)+2r_{k}}{p_{k}}=p_{k}-3+\frac{2r_{k}+2}{p_{k}}$. Thus, the decrease concerning $G_{k}$ is $1-2(r_{k}+1)\bigl(\frac{1}{p_{k}}-\frac{1}{p_{k}+1}\bigr)<1$, again a contradiction to (0). As a consequence, we have $p^{\prime}\leq p$ for $p_{k}=p^{\prime}$, and $p^{\prime}\leq p+1$ for $p_{k}=p^{\prime}-1$, hence verifying the stated upper bound.
If $q^{\prime}>0$, we show $p_{k}=p^{\prime}\leq p$, and that $p^{\prime}=p$ implies $q^{\prime}\leq q$. In this final case, copies of both $K_{p^{\prime}}$ and $K_{p^{\prime}+1}$ are present in the list $G_{1},\dots,G_{k-1}$. Referring to the proof of $d(\mathsf{L})\leq 1$ we see that $p_{k}<p^{\prime}$ would contradict (0) by moving an edge to $G_{k}$ from a copy of $K_{p^{\prime}+1}$, whereas $p_{k}>p^{\prime}$ would yield a similar contradiction by removing a vertex of degree $p_{k}$ from $G_{k}$ and extending a copy of $K_{p^{\prime}}$ to $K_{p^{\prime}+1}$. Thus, here we must have $p^{\prime}\leq p$, and the second claimed formula of the theorem applies. ∎
We note that the following side product of this result also follows:
  • If $0<r\leq\frac{1}{2}\,(p-1)$, then $M^{\mathsf{L}}(k,N)=M^{\mathsf{L}}(k,N-j)$ for $j=1,\dots,r$.
That is, in many cases it occurs that some (few or many) edges are unnecessary in the construction/verification of $M^{\mathsf{L}}(k,N)$.
Theorem 21.
Let $n\geq 3$ and $2\leq k\leq\binom{n}{2}$ be given, and choose the integer $p$ for which $k\binom{p}{2}\leq\binom{n}{2}<k\binom{p+1}{2}$ holds. Let $q,r$ be the nonnegative uniquely determined integers such that $\binom{n}{2}=k\binom{p}{2}+qp+r$, where $0\leq q<k$ and $0\leq r<p$. Then
[…]
Moreover, equality holds whenever $K_{n}$ admits an edge decomposition into
  • $q$ copies of $K_{p+1}$ and kqk-q copies of $K_{p}$, if $r=0$; or
  • $q$ copies of $K_{p+1}$, kq1k-q-1 copies of $K_{p}$, and one copy of any one graph $G\in\mathcal{F}_{p,\,r}$, if $0<r<p$.
Proof.
Since $M(k,n)\leq M^{\mathsf{L}}(k,\binom{n}{2})$, the claimed upper bound is a direct consequence of Theorem 20. The maximum sum for lists has been calculated from a combination of complete graphs $K_{p}$ and $K_{p+1}$, possibly together with a representative graph $G_{p,r}\in\mathcal{F}_{p,\,r}$. Thus, if $K_{n}$ admits an edge decomposition into the considered multiset of these graphs, then the list (multiset) also verifies that $M(k,n)$ attains $M^{\mathsf{L}}(k,\binom{n}{2})$. ∎
Concerning the possible cases of equality, the condition given in the theorem is sufficient but not necessary in general. This is shown by the following particular case.
Example 22.
Let $k=7$ and $n=8$. We have $e(K_{8})=\binom{8}{2}=28=12+12+4=2e(K_{4})+4e(K_{3})+e(G_{3,1})$, hence $M(7,8)\leq 2\,\mathrm{Mad}(K_{4})+4\,\mathrm{Mad}(K_{3})+\mathrm{Mad}(G_{3,1})=6+8+2=16$. Although two copies of $K_{4}$ do not pack with five copies of $K_{3}$ inside $K_{8}$, there exists a decomposition using the multiset $\{4*K_{3},2*(K_{4}-e),1*K_{4}\}$, which has $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})=4\cdot 2+2\cdot 2.5+3=16$. Indeed, Let $V(K_{8})=\{v_{1},\dots,v_{8}\}$. Embed $S(2,3,7)$ (the Fano plane) on the vertex set $V(K_{8})\diagdown\{v_{8}\}$ in a way that the triplets $(v_{1},v_{2},v_{3})$, $(v_{1},v_{4},v_{5})$, $(v_{1},v_{6},v_{7})$ are blocks of the STS. Extend these triplets with $v_{8}$, to obtain $K_{4}$ from the first and $K_{4}-e$ from the second and third (i.e., not using the edge $v_{1}v_{8}$ in the latter two). The other four blocks of the STS remain unchanged, yielding the required copies of $K_{3}$ for the decomposition of $K_{8}$.
It is worth noting that the above construction was possible because the transformation $(K_{4},G_{3,1})\longleftrightarrow(K_{4}-e,K_{4}-e)$, where $G_{3,1}$ is actually the paw graph, keeps the sum of $\mathrm{Mad}$ values unchanged (while the individual values do change).
Further non-uniform extremal constructions will be presented later. At this point we prove an asymptotic formula for $k$ growing proportionally to $\binom{n}{2}$.
5.2 The upper range
A Steiner system $S(2,p,n)$ is a collection of $p$-element subsets of an $n$-element underlying set, such that each pair of elements is contained in precisely one of those $p$-subsets. An $S(2,p,n)$ can equivalently be interpreted as a set of $\binom{n}{2}/\binom{p}{2}$ mutually edge-disjoint copies of $K_{p}$ in $K_{n}$. Beside $S(2,p,n)$ systems, there also exists a rich family of $\{K_{p},K_{p+1}\}$-decompositions of $K_{n}$, and more generally $\{K_{p},K_{p+1}\}$-packings. In this way, many exact formulas can be derived from Theorem 21. Below we explicitly formulate some of them.
Theorem 23.
Let $k=\binom{n}{2}-t$ with $0\leq t\leq\frac{1}{3}(n-1)^{2}$.
  • $(i)$
    If $t$ is even, then $M(k,n)=\binom{n}{2}-\frac{1}{2}t$, and the extremal multiset of subgraphs is $\{t/2*K_{3},\,(\,\binom{n}{2}-3t/2\,)*K_{2}\}$.
  • $(ii)$
    If $t$ is odd, then $M(k,n)=\binom{n}{2}-\frac{1}{2}(t+1)+\frac{1}{3}$, and the extremal multiset of subgraphs is $\{(t-1)/2*K_{3},\,(\,\binom{n}{2}-2-3(t-1)/2\,)*K_{2},\,1*P_{3}\}$.
Proof.
We apply Theorem 21. Using its notation, the assumed range of $t$ yields $p=2$, hence every optimal edge decomposition can be transformed to the claimed multiset of copies of single edges and triangles, together with one copy of $P_{3}$ if $t$ is odd. It is well known that such a packing of at least $\frac{1}{6}(n^{2}-2n-2)$ triangles exists indeed, for every $n$.
Suppose for a contradiction that there exists a further extremal multiset of graphs. Consider the very last step of the process of transformations described in the proof of Theorem 20. It is either $\{G_{1},G_{2}\}\rightarrow\{2*K_{3}\}$ or $\{G_{1},G_{2}\}\rightarrow\{K_{3},P_{3}\}$. Say, an edge $e$ is transposed from $G_{1}$ to $G_{2}$. Then $G_{1}-e\cong K_{3}$, hence $G_{1}\cong K_{3}$$\mathrm{leaf}$ or $G_{1}\cong K_{3}\cup K_{2}$. In either case, $e$ is a free edge in $G_{1}$. However, $K_{3}$ and $P_{3}$ have no free edges, which means that the insertion of $e$ caused an increase of $\mathrm{Mad}(G_{2})$. This contradicts the assumption that we started with an extremal multiset. ∎
Remark 24.
The limitation for $t$ in case of equality in Theorem 23 depends on the largest possible triangle packings in $K_{n}$. In the theory of block designs it was analyzed that the “leave graph” (formed by the edges of $K_{n}$ not covered by a largest partial Steiner triple system) is empty if $n\equiv 1$ or 33 $(\mathrm{mod}$ $6)$, $\frac{n}{2}K_{2}$ if $n\equiv 0$ or 22 $(\mathrm{mod}$ $6)$, $K_{1,3}\cup\frac{n-4}{2}K_{2}$ if $n\equiv 4$ $(\mathrm{mod}$ $6)$, or just the $4$-cycle $C_{4}$ if $n\equiv 5$ $(\mathrm{mod}$ $6)$. (See Theorem 3 of [22], or Table 40.22 in Section VI.40.4 of [11].) Accordingly, with a tight bound in Theorem 23, $t$ can be as large as $\frac{1}{3}n(n-1)$, $\frac{1}{3}n(n-2)$, $\frac{1}{3}(n-1)^{2}$, and $\frac{1}{3}(n^{2}-n-5)$, respectively; because in the last two cases, one $P_{3}$ can still be created optimally from $K_{1,3}$ or $C_{4}$.
6 Lower-bound constructions via Steiner systems and graph designs
6.1 Basic constructions
Theorem 25.
(Complete and partial Steiner systems.) Let $p\geq 3$.
  • $(i)$
    If $\binom{p}{2}\mid\binom{n}{2}$ and $p-1\mid n-1$, moreover if $n>n_{0}(p)$, then for $k=\frac{n(n-1)}{p(p-1)}$ we have $M(k,n)=n(n-1)/p=(p-1)\cdot k$.
  • $(ii)$
    As $n\to\infty$, for $k=\left\lfloor\frac{n(n-1)}{p(p-1)}\right\rfloor$ we have
    $\frac{n(n-1)}{p}-O(n)\leq M(k,n)\leq\frac{n(n-1)}{p}\,.$
Proof.
Under the given divisibility conditions of $(i)$, Wilson’s theorem [23] states that there exists an edge decomposition of $K_{n}$ into $\binom{n}{2}/\binom{p}{2}=k$ copies of $K_{p}$. Each copy has $\mathrm{Mad}=p-1$, implying $M(k,n)\geq n(n-1)/p$. This construction attains the maximum, due to Lemma 14 and Theorem 21.
For $(ii)$ assume that $n$ is large enough with respect to $p$, and choose the largest integer $n^{\prime}$ such that $n^{\prime}\leq n$ and $n^{\prime}\equiv 1$ or $p$ (mod p(p1))p(p-1)); then $0\leq n-n^{\prime}\leq(p-1)^{2}$. Observe that $n^{\prime}-1$ is a multiple of p1p-1, and $\binom{n^{\prime}}{2}$ is a multiple of $\binom{p}{2}$. Hence, due to $(i)$, for $k^{\prime}=\frac{n^{\prime}(n^{\prime}-1)}{p(p-1)}$ we have $M(k^{\prime},n^{\prime})=n^{\prime}(n^{\prime}-1)/p$. Here $k-k^{\prime}=\left\lfloor\frac{n(n-1)-n^{\prime}(n^{\prime}-1)}{p(p-1)}\right\rfloor$, which obviously is smaller than $\binom{n}{2}-\binom{n^{\prime}}{2}=\frac{n(n-1)-n^{\prime}(n^{\prime}-1)}{2}=(n-n^{\prime})\,\frac{n+n^{\prime}-1}{2}<\binom{p}{2}\cdot n$. From the $\binom{n}{2}-\binom{n^{\prime}}{2}$ edges of $K_{n}-K_{n^{\prime}}$ we form $k-k^{\prime}$ subgraphs of $K_{n}$ in an arbitrary way, and supplement them with a $K_{p}$-decomposition of $K_{n^{\prime}}$. Thus, we obtain $M(k,n)>M(k^{\prime},n^{\prime})=n^{\prime}(n^{\prime}-1)/p>n(n-1)/p-(p-1)\cdot n$. ∎
Theorem 26.
(Finite geometries.) Let $q$ be a prime power.
  • $(i)$
    If $n=k=q^{2}+q+1$, then $M(k,n)=(q^{2}+q+1)q=qk=\frac{2}{q+1}\binom{n}{2}$.
  • $(ii)$
    If $n=q^{2}$ and $k=q^{2}+q$, then $M(k,n)=(q+1)q(q-1)=(q-1)k=\frac{2}{q}\binom{n}{2}$.
  • $(iii)$
    If $n=q^{2}+q$ and $k=q^{2}+q+1$, then $M(k,n)=q^{3}+q^{2}-1=qk-q-1$.
  • $(iv)$
    If $n=q^{2}-1$ and $k=q^{2}+q$, then $M(k,n)=q^{3}-2q-1=(q-1)k-q-1$.
  • $(v)$
    If $n=q^{2}-q$ and $k=q^{2}+q-1$, then $M(k,n)=q^{3}-q^{2}-2q+1=(q-2)\cdot k+q-1$.
Proof.
Finite projective planes and in particular the projective Galois planes $PG(2,q)$, are Steiner systems $S(2,q+1,q^{2}+q+1)$, which are equivalent to decompositions of $K_{n}$ into copies of $K_{q+1}$. Similarly, finite affine planes, including the affine Galois planes $AG(2,q)$, are Steiner systems $S(2,q,q^{2})$, are equivalent to decompositions of $K_{n}$ into copies of $K_{q}$. In this way the formulas if $(i)$ and $(ii)$ follow by Theorem 25.
For the other parts we make the following modifications:
  • Concerning $(iii)$, pick a point of the projective plane $PG(2,q)$, and omit it from all the $q+1$ lines incident with it. This yields a set system over $q^{2}+q$ elements, with $q+1$ sets of size $q$ and $q^{2}$ sets of size $q+1$. The corresponding decomposition of $K_{q^{2}+q}$ has
    $M(q^{2}+q+1,q^{2}+q)\geq\sum_{i=1}^{k}\mathrm{Mad}(G_{i})=(q+1)(q-1)+q^{3}=q^{3}+q^{2}-1\,.$
  • Concerning $(iv)$, pick a point of the affine plane $AG(2,q)$, and omit it from all the $q+1$ lines incident with it. This yields a set system over $q^{2}-1$ elements, with $q+1$ sets of size q1q-1 and $q^{2}-1$ sets of size $q$. The corresponding decomposition of $K_{q^{2}-1}$ yields:
    $M(q^{2}+q,q^{2}-1)$ $\geq$ $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})$
    $=$ $(q+1)(q-2)+(q^{2}-1)(q-1)$
    $=$ $(q+1)(q^{2}-q-1)$
    $=$ $q^{3}-2q-1\,.$
  • Concerning $(v)$, pick a line $\ell$ of the affine plane $AG(2,q)$, and omit its points from all the $q^{2}$ lines which intersect $\ell$. This yields a set system over $q^{2}-q$ elements, with $q^{2}$ sets of size q1q-1 and q1q-1 sets of size $q$. (The q1q-1 lines parallel to $\ell$ remain intact.) The corresponding decomposition of $K_{q^{2}-q}$ has
    $M(q^{2}+q,q^{2}-1)$ $\geq$ $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\ =\ q^{2}\cdot(q-2)+(q-1)^{2}$
    $=$ $q^{3}-q^{2}-2q+1\ =\ (q-2)\cdot k+q-1\,.$
According to Theorem 25, all these three cases verify the exact value of $M(k,n)$, hence proving the formulas in (iii)(iii)$(v)$. ∎
The extremal structures in the cases of (iii)(iii)$(v)$ above are non-uniform. One can push this further, to obtain a larger class of tight constructions. For this, we apply particular types of $K$-designs. Given a set $K$ of positive integers, a pairwise balanced $K$-design ($K$-PBD, for short) of order $n$ and index $\lambda$ is a pair $(X,\mathcal{B})$, where $X$ is an underlying set of cardinality $n$ and $\mathcal{B}$ is a collection of blocks (subsets of $X$) such that each pair $\{x,x^{\prime}\}\subset X$ is contained in precisely $\lambda$ blocks. A parallel class is a set of mutually disjoint blocks whose union is $X$. A $K$-PBD $(X,\mathcal{B})$ is resolvable if $\mathcal{B}$ can be partitioned into parallel classes.
Theorem 27.
(Pairwise balanced K-designs with K = {p, p + 1}.)
$(i)$ Let $p\geq 3$ and $K=\{p,p+1\}$. Assume that $(X,\mathcal{B})$ is a $K$-PBD of order $n$ and index $1$, with all block sizes from $K$ (structurally equivalent to a $\{K_{p},K_{p+1}\}$-decomposition of $K_{n}$), with $b^{\prime}$ blocks of size $p$ and $b^{\prime\prime}$ blocks of size $p+1$. Then, for $k=b^{\prime}+b^{\prime\prime}$, we have $M(k,n)=(p-1)\cdot k+b^{\prime\prime}$.
$(ii)$ Let $p\geq 3$ and $n\equiv p$ ((mod p(p1))p(p-1)). Assume that at least one of the following conditions holds:
  • $n>\exp\{\exp\{p^{12p^{2}}\}\}$ and $n^{\prime}\leq\frac{n-1}{p-1}$;
  • $n^{\prime}<10^{-4}n/p^{5/2}$.
Then the following holds.
  • $(a)$
    If there exists an $S(2,p,n^{\prime})$-system, then for $k=\frac{n(n-1)}{p(p-1)}+\frac{n^{\prime}(n^{\prime}-1)}{p(p-1)}$ we have $M(k,n+n^{\prime})=(p-1)\cdot k+(n\cdot n^{\prime})/p$.
  • $(b)$
    If there exists an $S(2,p+1,n^{\prime})$-system, then for $k=\frac{n(n-1)}{p(p-1)}+\frac{n^{\prime}(n^{\prime}-1)}{p(p+1)}$ we have $M(k,n+n^{\prime})=(p-1)\cdot k+(n\cdot n^{\prime})/p+n^{\prime}(n^{\prime}-1)/(p+1)$.
Proof.
Similarly as above, Theorem 25 implies that every decomposition of $K_{n}$ into any $b^{\prime}$ copies of $K_{p}$ and $b^{\prime\prime}$ copies of $K_{p+1}$ provides the exact value of $M(b^{\prime}+b^{\prime\prime},n)$. This verifies $(i)$.
In order to prove $(ii)$ we need to show that the corresponding decompositions of $K_{n+n^{\prime}}$ exist under the conditions imposed on $n$, $n^{\prime}$, and $p$. We view $K_{n+n^{\prime}}$ as the complete join $K_{n}+K_{n^{\prime}}$.
For the case of $n^{\prime}\leq\frac{n-1}{p-1}$ and $n>\exp\{\exp\{p^{12p^{2}}\}\}$, Chang [9, Theorem 7.2] proved that there exists a resolvable $S(2,p,n)$ system. Using its blocks as vertex sets of complete subgraphs in $K_{n}$, from the parallel classes of a resolution we obtain a decomposition of $K_{n}$ into $\frac{n-1}{p-1}$ copies of $\frac{n}{p}K_{p}$.
Each of the $n^{\prime}$ vertices in $K_{n^{\prime}}$ can be used to extend the $n/p$ copies of $K_{p}$ in a parallel class to the same number of $K_{p+1}$ subgraphs in $K_{n+n^{\prime}}$, distinct classes taken for distinct vertices. Note that these cover all the edges connecting $V(K_{n})$ to $V(K_{n^{\prime}})$. The other $\frac{n-1}{p-1}-n^{\prime}$ parallel classes and the blocks of the assumed $S(2,p,n^{\prime})$ or $S(2,p+1,n^{\prime})$ system ensure that a required decomposition of $K_{n+n^{\prime}}$ into $k$ complete subgraphs (each of them being $K_{p}$ or $K_{p+1}$) exists, indeed.
For the case of $n^{\prime}$ much smaller than $n/p$ we need not assume a huge $n$; namely, the condition $p\mid n$ will suffice. Also here we shall need $n^{\prime}$ parallel classes $\frac{n}{p}K_{p}$, which we find using the Hajnal–Szemerédi Theorem [17]. It states that, for every positive integer $s$, every graph with maximum degree at most $s$ has an equitable coloring with $s+1$ colors. Putting s=n/p1s=n/p-1, equitable means that all color classes have size $p$; those classes will be vertex sets of the copies of $K_{p}$, together forming parallel classes $F_{i}\cong\frac{n}{p}K_{p}$. We select $F_{1},F_{2},\dots,F_{n^{\prime}}$ sequentially. The first one, $F_{1}$, is just any choice of $\frac{n}{p}K_{p}$. Let us denote $E_{i}=E(F_{i})$. Having chosen $F_{1},\dots,F_{j}$, consider the graph whose edge set is $E_{1}\cup\cdots\cup E_{j}$. This graph is regular of degree (p1)j<n/p1(p-1)j<n/p-1, therefore it admits an equipartite coloring with $n/p$ colors. This coloring yields $F_{j+1}$. The procedure leads to $n^{\prime}$ parallel classes as needed. The edge set of each $F_{i}$ is supplemented with a distinct vertex of $K_{n^{\prime}}$, hence yielding $(n\cdot n^{\prime})/p$ copies of $K_{p+1}$. We can also use the blocks of the assumed $S(2,p,n^{\prime})$ or $S(2,p+1,n^{\prime})$ to obtain further complete subgraphs $K_{p}$ or $K_{p+1}$ in $K_{n+n^{\prime}}$.
It remains to decompose the graph $H$ having the edge set $E(H)=E(K_{n})-\cup_{i=1}^{n^{\prime}}E_{i}$ into copies of $K_{p}$. Gustavsson [16] proved that this can be done whenever the following conditions are satisfied: $\binom{p}{2}\mid e(H)$, the degree of every vertex is a multiple of p1p-1, and $\delta(H)>(1-\epsilon)n$, where $\epsilon=\epsilon(p)>0$ is a sufficiently small absolute constant. Making a substantial improvement on $\epsilon$, Glock et al. [14, Corollary 1.6] proved that $\epsilon\geq 10^{-4}p^{-3/2}$ is a suitable choice for $K_{p}$-decomposition. Thus, the described strategy is feasible whenever at most $\epsilon n/(p-1)$ parallel classes are needed. This completes the proof of the theorem. ∎
Remark 28.
The principle of proof for part $(ii)$ also yields the following more general recursion:
  • $(iii)$
[…]
    $M(k,n+n^{\prime})=(p-1)\cdot k+b_{n}^{\prime\prime}+b_{n^{\prime}}^{\prime\prime}+(n\cdot n^{\prime})/p\,.$
It was shown in Theorem 23 that the behavior of $M(k,n)$ for $k\geq\frac{1}{3}\binom{n}{2}$ is understood fairly well. The next result gives some information concerning the remaining part of the quadratic range of $k$.
Theorem 29.
(Fixed proportional growth of k.) Let $r\geq 3$ be any rational number. There exist infinitely many $n$ such that $M(k,n)$ for $k=\frac{1}{r}\binom{n}{2}$ is determined by a $\{K_{p},K_{p+1}\}$-design, where $p\geq 3$ is the integer satisfying $\binom{p}{2}\leq r<\binom{p+1}{2}$.
Proof.
Let us note first that the case $r=\binom{p}{2}$ is already settled in part $(i)$ of Theorem 25. Therefore, in the sequel we assume $\binom{p}{2}<r<\binom{p+1}{2}$.
We search for an edge decomposition of $K_{n}$ into the members of the multiset $\{xk*K_{p}\,,(1-x)k*K_{p+1}\}$, with a suitable choice of $x\in(0,1)$. Since $e(K_{n})=\binom{n}{2}=rk$, this requires
$xk\binom{p}{2}+(1-x)k\binom{p+1}{2}=rk\,$,
thus
$x=\frac{p+1}{2}-\frac{r}{p}\,.$
Since the integer $p$ is determnimned by $r$, we see that $x$ is a function of $r$, a rational number between 0 and 1 because $\frac{p-1}{2}<\frac{r}{p}<\frac{p+1}{2}$ by the choice of $p$. In particular, $x$ is independent of the actual choice of the number $n$ of vertices.
Assume $x=s/t$ ($s,t\in\mathbb{N}$ ,  $1\leq s\leq t-1$), and let $H=sK_{p}\cup(t-s)K_{p+1}$. Then $\delta(H)=p$ and $\Delta(H)=p+1$. Thus, Wilson’s theorem implies that $K_{n}$ admits an $H$-decomposition whenever $\binom{n}{2}$ is a multiple of $e(H)$ and $n$ is sufficiently large. Here $e(H)=rt$, hence $K_{n}$ is decomposed into $\binom{n}{2}/(rt)$ copies of $H$, which then further decompose into $\frac{s}{rt}\binom{n}{2}$ copies of $K_{p}$ and $\frac{t-s}{rt}\binom{n}{2}$ copies of $K_{p+1}$. Consequently, a $\{K_{p},K_{p+1}\}$-design is obtained with exactly $\frac{1}{r}\binom{n}{2}$ subgraphs. By the “moreover” part of Theorem 21 we obtain:
$M(k,n)=\left(\frac{s\cdot(p-1)}{rt}+\frac{(t-s)\cdot p}{rt}\right)\!\binom{n}{2}=\frac{tp-s}{rt}\binom{n}{2}$,
implying the assertion. ∎
In particular, we obtain:
Corollary 30.
For any rational $r\geq 1$ there are infinitely many $n$ such that $M(k,n)$ is an integer for $k=\frac{1}{r}\binom{n}{2}$.
Proof.
Assuming $r=a/b$, apply the formula at the end of the proof of Theorem 29, and restrict the solutions to those $n$ for which $\binom{n}{2}$ is a multiple of $ta$. ∎
6.2 Constructions via substitution and the asymptotic growth of M (k , n)
Here we first consider the cases of triangular numbers $k=\binom{t+1}{2}$. Since $M(k,n)<\sqrt{k}\,n$ by Corollary 17, the following lower bound is tight apart from a multiplicative constant in the entire range of $k$. Substantial improvement for large $k$ will be given later in this section.
Proposition 31.
Let $k=\binom{t+1}{2}$ for some $t\geq 2$, and assume $t\mid n$. Then $M(k,n)\geq\frac{t+1}{2}\,n-t>\sqrt{k/2}\,(n-2)$.
Proof.
Partition $V(K_{n})$ into $t$ sets $V_{1},\dots,V_{t}$ of cardinality $n/t$ each. Define $t+\binom{t}{2}$ subgraphs indexed as $G_{i}$ for $1\leq i\leq t$ and $G_{i,j}$ for $1\leq i<j\leq t$, as follows. Each $G_{i}$ is a copy of $K_{n/t}$ with $V(G_{i})=V_{i}$, and each $G_{i,j}$ is a copy of $K_{\frac{n}{t},\frac{n}{t}}$ with $V(G_{i,j})=V_{i}\cup V_{j}$. All these graphs are regular. More explicitly, we have obtained $t$ graphs of degree n/t1n/t-1, and $\frac{1}{2}t(t-1)$ graphs of degree $n/t$. Consequently, $M(k,n)\geq n-t+\frac{1}{2}(t-1)n$. The last inequality is valid because $k<(t+1)^{2}\!/2$. ∎
Theorem 32.
Let $v>r>2$ be such that a Steiner system $S(2,r,v)$ exists. If $k=\frac{v(v-1)}{r(r-1)}$, and $n$ is any integer divisible by $v$, then $M(k,n)\geq\frac{v}{r}\left(n-1\right)$.
Proof.
Let $(X,\mathcal{B})$ be an $S(2,r,v)$ with point set $X$ and blocks $B_{1},\dots,B_{k}$. Replace the points $x\in X$ with mutually disjoint sets $V_{x}$ of cardinality $n/v$. Then, viewing $\cup_{x\in X}V_{x}$ as the vertex set of $K_{n}$, we obtain $k$ sets $Y_{i}:=\cup_{x\in B_{i}}V_{x}$ of cardinality $rn/v$. For $i=1,\dots,k$ let $H_{i}$ be the complete $r$-partite graph whose edges are the vertex pairs $y,y^{\prime}$ where $y,y^{\prime}\in Y_{i}$, say $y\in V_{x}$ and $y^{\prime}\in V_{x^{\prime}}$, such that $x\neq x^{\prime}$. Distribute the remaining edges of $K_{n}$ (i.e., those which are contained in some $V_{x}$) arbitrarily among the $H_{i}$. In this way we obtain an edge decomposition of $K_{n}$ into $k$ subgraphs; denote them by $G_{1},\dots,G_{k}$.
The average number of edges in the $G_{i}$ is $\frac{n(n-1)}{2k}$, and each of them has order $rn/v$; hence the average of their $d(G_{i})$ is equal to $\frac{v}{rk}\left(n-1\right)$. Thus, $M(k,n)\geq\sum_{i=1}^{k}\mathrm{Mad}(G_{i})\geq\sum_{i=1}^{k}d(G_{i})=\frac{v}{r}\left(n-1\right)$. ∎
Corollary 33.
Let $q$ be a prime power.
  • $(i)$
    If $k=q^{2}+q+1$ and $k\mid n$, then $M(k,n)\geq(q+\frac{1}{q+1})(n-1)$.
  • $(ii)$
    If $k=q^{2}+q$ and $q^{2}\mid n$, then $M(k,n)\geq qn-q$.
Proof.
Apply Theorem 32 to the projective planes $S(2,q+1,q^{2}+q+1)$ and affine planes $S(2,q,q^{2})$, with $v=q^{2}+q+1$ or $v=q^{2}$ and $r=q+1$ or $r=q$, respectively. ∎
Now we are in a position to prove that $M(k,n)$ as a function of $n$ grows essentially with $\sqrt{k}\,n$.
Theorem 34.
For every $n$ and every fixed $k$ we have
$(1-\epsilon_{k})\sqrt{k}\,n<M(k,n)<\sqrt{k}\,n\,$,
where $\epsilon_{k}\to 0$ as $k\to\infty$. In convergence,
$\lim_{k\to\infty}\Bigl(\liminf_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}\Bigr)=\lim_{k\to\infty}\Bigl(\limsup_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}\Bigr)=1\,.$
Proof.
Recall that the upper bound $\sqrt{k}\,n$ is valid by Corollary 17. Moreover, if $k=q^{2}+q+1$ where $q$ is a prime power and $n$ is a multiple of $k$, then the lover bound holds in a stronger form. Namely, we can write
$(\sqrt{k}-1/2)\,n-c_{k}$
instead of $(1-\epsilon_{k})\sqrt{k}\,n$, choosing $q+\frac{1}{q+1}$ (which is less than $\sqrt{k}$ ) for the constant $c_{k}$. Indeed, from part $(i)$ of Corollary 33 we obtain
$M(k,n)+q+\frac{1}{q+1}\geq(q+\frac{1}{q+1})\,n>(\sqrt{q^{2}+q+1}-1/2)\,n=(\sqrt{k}-1/2)\,n$
where the inequality in the middle can easily be checked.
For orders not divisible by $k$, let us consider $n^{\prime}=n+t=ks+t$, where $1\leq t\leq k-1$. By monotonicity with respect to $n$ (due to Proposition 2) we know
$M(k,n^{\prime})$ $\geq$ $M(k,n)\ >\ (\sqrt{k}-1/2)\,(n^{\prime}-t)-\sqrt{k}$
$\geq$ $(\sqrt{k}-1/2)\,n^{\prime}-\sqrt{k}-(k-1)(\sqrt{k}-1/2)$
$=$ $(\sqrt{k}-1/2)\,n^{\prime}-k^{3/2}+k/2-1/2\,.$
Hence $k^{3/2}$ is a suitable choice for $c_{k}$ whenever $k=q^{2}+q+1$ with any prime power $q$.
Suppose that $k$ cannot be written in this form. Let then $q$ be the smallest prime power such that $q^{2}+q+1>k$. It is well known that there exists a sublinear function $f(x)=o(x)$ such that the interval [xf(x),x][x-f(x),x] contains a prime number for every $x>2$. (The current record concerning the growth of $f$ seems to be $o(x^{0.525})$, see [3].) So, we can choose a prime (or prime power) $q^{\prime}$ with $q-q^{\prime}=o(q)$. Let $k^{\prime}=(q^{\prime})^{2}+q^{\prime}+1$; certainly $k^{\prime}<k$, moreover $k-k^{\prime}=o(k)$ and $\sqrt{k}-\sqrt{k^{\prime}}=o(\sqrt{k}\,)$ as $k$ grows. Since $M(k,n)$ is monotone increasing in $k$, again by Proposition 2, we obtain:
$M(k,n)>M(k^{\prime},n)>(\sqrt{k^{\prime}}-1/2)\,n-c_{k^{\prime}}\geq(\sqrt{k}-\epsilon_{k}\sqrt{k}\,)\,n$
with a suitably chosen $\epsilon_{k}>0$.
For the limits $s_{k}:=\liminf_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}$ and $t_{k}:=\limsup_{n\to\infty}\frac{M(k,n)}{\sqrt{k}\,n}$ this yields $1-o_{k}(1)\leq s_{k}\leq t_{k}\leq 1$, implying the theorem. ∎
In the last inequality above, the constant $\epsilon_{k}$ can be estimated via the $f(x)=o(x^{0.525})$ theorem quoted in the proof. The extension of the theorem for non-fixed $k$ faces the difficulty that the $o$-bound on $f$ requires $x$ to be large, but $n$ puts limitations on that.
Part $(i)$ of Corollary 33 can be extended also for values of $k$ which are not of the form $q^{2}+q+1$. This offers a replacement of the above error term $\epsilon_{k}\sqrt{k}\,n$ with an explicit function.
Proposition 35.
Let $k=q^{2}+q+1+r$, where $q$ is a prime power and $1\leq r\leq q^{2}+q+1$. If $(q^{2}+q+1)|n$, then $M(k,n)\geq(q+\frac{1}{q+1})\,n+\frac{rq}{(q+1)(q^{2}+q+1)}\,n-(q+r)$.
Proof.
For this purpose we need to specify the construction of Theorem 32 more strictly. We apply the fact that every projective Galois plane has a cyclic representation (e.g., via difference sets). Let $x_{1},\dots,x_{q^{2}+q+1}$ and $\ell_{1},\dots,\ell_{q^{2}+q+1}$ be the points and lines, respectively, of $PG(2,q)$, such that $x_{i}\in\ell_{i}$ for $i=1,\dots,q^{2}+q+1$ and the rotation $x_{i}\mapsto x_{i+1}$ performed modulo $q^{2}+q+1$ is an automorphism of $PG(2,q)$. Replace each $x_{i}$ with a set $X_{i}$ of cardinality $n/(q^{2}+q+1)$, creating a complete equipartite graph $H_{i}$ from $\ell_{i}$ with $q+1$ vertex classes.
If $r+1\leq i\leq q^{2}+q+1$, we let $G_{i}$ be the graph obtained from $H_{i}$ by inserting the edges of the complete graph whose vertex set is $X_{i}$. This $G_{i}$ has $\frac{1}{q+1}|G_{i}|$ vertices of degree $|G_{i}|-1$ (namely the vertices in $X_{i}$), and its other $(1-\frac{1}{q+1})|G_{i}|$ vertices have degree $(1-\frac{1}{q+1})|G_{i}|$. Recalling that $|G_{i}|=\frac{q+1}{q^{2}+q+1}\,n$, we obtain
$\sum_{i=r+1}^{q^{2}+q+1}d(G_{i})$ $=$ $\sum_{i=r+1}^{q^{2}+q+1}\frac{1}{|G_{i}|}\left(\frac{1}{q+1}|G_{i}|\left(|G_{i}|-1\right)+\left(1-\frac{1}{q+1}\right)^{2}|G_{i}|^{2}\right)$ (1)
$=$ $\sum_{i=r+1}^{q^{2}+q+1}\left(\left(\frac{1}{q+1}+\left(\frac{q}{q+1}\right)^{2}\right)|G_{i}|-\frac{1}{q+1}\right)$
$=$ $\left(q^{2}+q+1-r\right)\left(\frac{q^{2}+q+1}{(q+1)^{2}}\left(\frac{q+1}{q^{2}+q+1}\,n\right)-\frac{1}{q+1}\right)$
$\geq$ $\frac{q^{2}+q+1-r}{q+1}\,n-q\,.$
If $1\leq i\leq r$, we simply put $G_{i}=H_{i}$. These graphs are regular of degree $(1-\frac{1}{q+1})|G_{i}|$, hence
$\sum_{i=1}^{r}d(G_{i})$ $=$ $r\left(\left(1-\frac{1}{q+1}\right)|G_{i}|\right)$ (2)
$=$ $\frac{rq}{q+1}\,\frac{q+1}{q^{2}+q+1}\,n\,.$
Finally, if $q^{2}+q+2\leq i\leq q^{2}+q+1+r$, we let $G_{i}$ be the complete graph on the vertex set $X_{i}$. These $r$ graphs are regular of degree $|X_{i}|-1$, hence
$\sum_{i=q^{2}+q+2}^{q^{2}+q+1+r}d(G_{i})$ $=$ $\frac{r}{q^{2}+q+1}\,n-r.$ (3)
Summing up (1)+(2)+(3) we obtain
$M(k,n)$ $\geq$ $\sum_{i=1}^{k}\mathrm{Mad}(G_{i})$
$=$ $\left(q+\frac{1}{q+1}\right)\,n+\left(\frac{q+1}{q^{2}+q+1}-\frac{1}{q+1}\right)\,rn-q-r\,$,
proving the assertion. ∎
An explicit lower bound of similar flavor can also be given for the cases where $k$ is big, but not much bigger than the number of blocks in a Steiner system.
Proposition 36.
Assume that $k=\binom{n}{2}/\binom{p}{2}$, and that a Steiner system $S(2,p,n)$ exists. If $1\leq r\leq 2k$, then $M(k+r,n)\geq M(k,n)+rp/4-cr$ for a constant $c$.
Proof.
We know from Theorem 21 that $M(k,n)$ for $k=\binom{n}{2}/\binom{p}{2}$ is generated via $S(2,p,n)$ whenever such a Steiner system exists. In that construction, every graph $G_{i}$ in the decomposition of $K_{n}$ is a copy of $K_{p}$. Applying Theorem 18 and the first subcase of Proposition 19, we decompose selected copies of $K_{p}$ into two subgraphs $H_{i,1},H_{i,2}$ with $\mathrm{Mad}(H_{i,1})+\mathrm{Mad}(H_{i,2})\geq 5p/4-3/2$ or into three subgraphs $H_{i,1},H_{i,2},H_{i,3}$ with $\mathrm{Mad}(H_{i,1})+\mathrm{Mad}(H_{i,2})+\mathrm{Mad}(H_{i,3})\geq 3p/2-c_{3}$. Increasing the number of subgraphs with $r$, this operation increases $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$ with at least rp/4crrp/4-cr, where $c$ is relatively small. ∎
Also, the following recursive lower bound can be stated.
Proposition 37.
For every $a,b,t\in\mathbb{N}$ with $b\geq 3$, $t\geq 2$ and $1\leq a\leq\binom{b}{2}$ we have $M(a+b,tb)\geq t\cdot(M(a,b)+b)-b$.
Proof.
Applying the principle of the preceding constructions, replace each vertex $x_{i}$ of a structure verifíying $M(a,b)$ with a set $X_{i}$ of cardinality $t$, blow up the corresponding graphs $G_{1},\dots,G_{a}$ accordingly (hence multiplying their $\mathrm{Mad}$ values by $t$), and create $b$ further graphs $G_{a+1},\dots,G_{a+b}$ regular of degree t1t-1, $G_{a+i}$ being the complete graph on the vertex set $X_{i}$ for $i=1,\dots,b$. ∎
Applying the above implication for $t=n/b$, we may abbreviate the conclusion as $M(a,b)\Longrightarrow M(a+b,n)\gtrsim c_{a,b}\cdot n$ for an asymptotic lower bound as $n$ gets large, where $c_{a,b}$ abbreviates the value $\frac{1}{b}\cdot M(a,b)+1$. Some pairs $a,b$ provide lower bounds that coincide with those given in Proposition 19:
  • $M(1,2)\Longrightarrow M(3,n)\gtrsim 3n/2$;
  • $M(1,3)\Longrightarrow M(4,n)\gtrsim 5n/3$;
  • $M(3,3)\Longrightarrow M(6,n)\gtrsim 2n$.
In the same way also $M(k,n)\Longrightarrow M(k+n,tn)\geq t\cdot M(k,n)+(t-1)\cdot n$ can be obtained.
7 Applications to other graph invariants
In this section we derive implications of the above results for other graph invariants, namely:
  • clique number, $\omega(G)$ (largest number of mutually adjacent vertices in $G$);
  • chromatic number, $\chi(G)$ (minimum number of colors in a vertex coloring such that no two adjacent vertices have the same color);
  • choice number or list chromatic number, $\mathrm{ch}(G)$ (minimum size $m$ of color lists $L_{v}$ assigned to the vertices $v\in V(G)$ guaranteeing that there exists a coloring $\varphi:V(G)\to\cup_{v\in V(G)}L_{v}$ with $\varphi(v)\in L_{v}$ for all $v\in V(G)$ and $\varphi(v)\neq\varphi(w)$ for all $vw\in E(G)$ whenever $|L_{v}|=m$ for all $v\in V(G)$);
  • degeneracy, $\delta{}^{\mathsf{+}}(G)$ (smallest integer $d$ such that $G$ admits a vertex order $v_{1},\dots,v_{n}$ in which each $v_{j}$ has at most $d$ preceding neighbors $v_{i}$ with $i<j$, equal to $\max\{\delta(H)\}$ taken over all induced subgraphs $H\subseteq G$);
  • coloring number, $\mathrm{col}(G)=\delta{}^{\mathsf{+}}(G)+1$, also called Szekeres–Wilf number;
  • highest induced vertex connection, $\kappa{}^{\mathsf{+}}(G)$ (maximum vertex connection number $\kappa(H)$ over all induced subgraphs $H\subseteq G$);
  • highest induced edge connection, $\lambda{}^{\mathsf{+}}(G)$ (maximum edge connection number $\lambda(H)$ over all induced subgraphs $H\subseteq G$).
Applying the notational convention described in the introduction, we write $\omega(k,n)$, $\chi(k,n)$, $\delta{}^{\mathsf{+}}(k,n)$, $\mathrm{col}(k,n)$, $\kappa{}^{\mathsf{+}}(k,n)$, $\lambda{}^{\mathsf{+}}(k,n)$ for the largest possible sum $\sum_{j=1}^{k}{\mathfrak{p}}(G_{j})$ taken over all edge decompositions of $K_{n}$ into $k$ subgraphs $G_{1},\dots,G_{k}$, where ${\mathfrak{p}}\in\{\omega,\chi,\delta{}^{\mathsf{+}},\mathrm{col},\kappa{}^{\mathsf{+}},\lambda{}^{\mathsf{+}}\}$.
Since the inequality chains $\mathrm{Mad}(G)+1\geq\mathrm{col}(G)\geq\mathrm{ch}(G)\geq\chi(G)\geq\omega(G)$ and $\mathrm{Mad}(G)\geq\delta{}^{\mathsf{+}}(G)\geq\lambda{}^{\mathsf{+}}(G)\geq\kappa{}^{\mathsf{+}}(G)$ are valid for every graph $G$, also the corresponding extremal functions satisfy the following numerical relations for all $n\geq 3$ and $2\leq k\leq\binom{n}{2}$ :
$\omega(k,n)\leq\chi(k,n)\leq\mathrm{ch}(k,n)\leq\mathrm{col}(k,n)\leq M(k,n)+k\,$,
$\kappa{}^{\mathsf{+}}(k,n)\leq\lambda{}^{\mathsf{+}}(k,n)\leq\delta{}^{\mathsf{+}}(k,n)\leq M(k,n)\,.$
Bickle [5] conjectures that $\delta{}^{\mathsf{+}}(k,n)=\kappa{}^{\mathsf{+}}(k,n)=\lambda{}^{\mathsf{+}}(k,n)$ holds for all $k,n$. We shall see in Theorem 40 that this is indeed the case at least when a supplementary assumption is imposed.
Füredi et al. studied four of the above seven parameters; next we summarize their findings.
Theorem 38 (Füredi, Kostochka, Stiebitz, Škrekovski, West [13]).
  1. 1.
    $\omega(k,n)\leq n+\binom{k}{2}$, with equality for all $n\geq\binom{k}{2}$.
  2. 2.
    $\chi(k,n)\leq n+7^{k}$.
  3. 3.
    $\mathrm{ch}(k,n)\leq n+3k!\sqrt{1+8n\ln n}$, moreover $\mathrm{ch}(k,n)\geq n+ck\ln(n/\!\sqrt{k}\,)$ for a constant $c>0$ if $k$ is a triangular number, $k=\binom{t+1}{2}$, and $t\mid n$.
  4. 4.
    $\mathrm{col}(k,n)\leq\sqrt{k}\,n+k$, moreover $\mathrm{col}(k,n)\geq(\sqrt{k}-1)n+k$ if $k=q^{2}+q+1$ for a prime power $q$, and $k\mid n$.
  5. 5.
    $\mathrm{col}(k,n)\leq(k-1)(n+1)/2$ if $k\geq 5$; for smaller $k$, $\mathrm{col}(3,n)=(3n+3)/2$ and $\mathrm{col}(4,n)=(5n+7)/3$.
Let us recall further the following Norhaus–Gaddum-type result (converted to the present notation) from the paper which initiated the study of choice number.
Theorem 39 (Erdős, Rubin, Taylor [12]).
For every $n\in\mathbb{N}$, $\mathrm{ch}(2,n)\leq\mathrm{col}(2,n)\leq n+1$.
Comparing Theorem 39 with Theorem 18 it turns out that $M(2,n)\sim\frac{5}{4}\,n$ is much bigger than $\mathrm{col}(2,n)$. For $k=3$ and $k=4$ it is not clear whether the situation between $M(k,n)$ and $\mathrm{col}(k,m)$ is similar, as the corresponding constructions in Proposition 19 yield lower bounds which coincide with those in Part 5 of Theorem 38. However, as $k$ gets large and $n\to\infty$, the relative gap between $\mathrm{col}(k,n)$ and $M(k,n)$ gets small. What is more, our results on $M(k,n)$ improve substantially the earlier estimates on several functions ${\mathfrak{p}}(k,n)$ in certain ranges of the parameters.
Let us recall that the list version $M^{\mathsf{L}}(k,\binom{n}{2})$ of $M(k,n)$ (and more generally, $M^{\mathsf{L}}(k,N)$ for all $N\geq k$) was determined exactly for all $n\geq 3$ and all $2\leq k\leq\binom{n}{2}$ in Theorem 20. It has also been proved that every list of graphs being extremal for $M^{\mathsf{L}}(k,\binom{n}{2})$ can be transformed to one of two types:
  • Type 1: $q$ copies of $K_{p+1}$ and kqk-q copies of $K_{p}$, if $r=0$;
  • Type 2: $q$ copies of $K_{p+1}$, kq1k-q-1 copies of $K_{p}$, and one copy of any one graph $G\in\mathcal{F}_{p,\,r}$, if $0<r<p$;
where $\binom{n}{2}=k\binom{p}{2}+qp+r$, with $0\leq q<k$ and $0\leq r<p$. As stated in Theorem 21, we have $M(k,n)=M^{\mathsf{L}}(k,\binom{n}{2})$ whenever $M(k,n)$ is attained by a packing of Type 1 or Type 2. Moreover, a bunch of results above show that this situation occurs for infinitely many pairs $k,n$. Consequences can be stated for other parameters as well.
Theorem 40.
If $M(k,n)$ is attained by a packing of Type 1 or Type 2, and in case of Type 2 with $r\geq(p-1)/2$ the packing is established using a graph $G\in\mathcal{G}_{p,r}$ with $\omega(G)=p$, then
$\omega(k,n)=\chi(k,n)=\mathrm{ch}(k,n)=\mathrm{col}(k,n)=\left\lfloor M(k,n)\right\rfloor+k\,$,
and
$\kappa{}^{\mathsf{+}}(k,n)=\lambda{}^{\mathsf{+}}(k,n)=\delta{}^{\mathsf{+}}(k,n)=\left\lfloor M(k,n)\right\rfloor\,.$
Proof.
In a packing of Type 1, all graphs are complete, and we have $\kappa{}^{\mathsf{+}}(K_{p})=\kappa(K_{p})=p-1=\mathrm{Mad}(K_{p})$ and $\omega(K_{p})=p=\mathrm{Mad}(K_{p})+1$ (and of course the values for $K_{p+1}$ are larger by eaxctly 1). In case of Type 2, with $G\in\mathcal{G}_{p,r}$ being the single non-complete graph in the packing (where $0<r<p$), we have $p-1<\mathrm{Mad}(G)<p$. Then $\kappa{}^{\mathsf{+}}(G)=\left\lfloor\mathrm{Mad}(G)\right\rfloor$ and $\omega(G)=\left\lfloor\mathrm{Mad}(G)\right\rfloor+1$. Summing up for $G_{1},\dots,G_{k}$, the theorem follows. ∎
As a matter of fact, the behavior of several functions is very similar.
Theorem 41.
For every parameter ${\mathfrak{p}}\in\{\kappa{}^{\mathsf{+}},\lambda{}^{\mathsf{+}},\delta{}^{\mathsf{+}}\}$ we have
$(1-\epsilon_{k})\sqrt{k}\,n\leq{\mathfrak{p}}(k,n)\leq M(k,n)<\sqrt{k}\,n$
for every fixed $k$, where $\epsilon_{k}\to 0$ as $k\to\infty$.
Proof.
The upper bound $\sqrt{k}$ is valid by Corollary 17. Since $\mathrm{col}(G)-1=\delta{}^{\mathsf{+}}(G)\geq\lambda{}^{\mathsf{+}}(G)\geq\kappa{}^{\mathsf{+}}(G)$ holds for every graph $G$, it remains to prove the lower bound $\kappa{}^{\mathsf{+}}(k,n)\geq(1-\epsilon_{k})\sqrt{k}\,n$.
For this, we apply the construction by which we derived the asymptotic estimates of Theorem 34 along similar lines. Let $k^{\prime}\leq k$ be the largest integer of the form $k^{\prime}=q^{2}+q+1$, where $q$ is a prime power, and $n^{\prime}$ be the largest multiple of $k^{\prime}$ with $n^{\prime}\leq n$. Then $K_{n^{\prime}}$ admits a packing $G_{1},\dots,G_{k^{\prime}}$ of $k^{\prime}$ edge-disjoint copies of the complete $(q+1)$-partite graph $K_{\frac{n^{\prime}}{k^{\prime}},\frac{n^{\prime}}{k^{\prime}},\dots,\frac{n^{\prime}}{k^{\prime}}}$. If fewer than $\frac{q}{q^{2}+q+1}\,n^{\prime}$ vertices are removed from it, then more than one vertex class contains a non-deleted vertex. Picking two of those vertices from distinct classes, they are adjacent and dominate the rest of $K_{\frac{n^{\prime}}{k^{\prime}},\frac{n^{\prime}}{k^{\prime}},\dots,\frac{n^{\prime}}{k^{\prime}}}$. Thus, $\kappa{}^{\mathsf{+}}(G_{i})\geq\frac{q}{q^{2}+q+1}\,n^{\prime}$ holds for all $q\leq i\leq k^{\prime}$.
Distribute the edges of $E(K_{n})\diagdown(\cup_{i=1}^{k^{\prime}}E(G_{i}))$ among the current graphs $G_{i}$ and the further $k-k^{\prime}$ graphs $G_{k^{\prime}+1},\dots,G_{k}$ to be constructed. Using the facts $k-k^{\prime}=o(k)$ and $n-n^{\prime}<k$, we obtain
$\kappa{}^{\mathsf{+}}(k,n) \displaystyle\geq$ $\sum_{i=1}^{k^{\prime}}\kappa{}^{\mathsf{+}}(G_{i})\ \geq\ \frac{q}{q^{2}+q+1}\,k^{\prime}n^{\prime}\ =\ qn^{\prime}$
$>$ $(\sqrt{k^{\prime}}-1)\,n^{\prime}\ \geq\ (1-\epsilon_{k})\sqrt{k}\,n\,.$
From the first three parts of Theorem 38 one can see that the behavior of ${\mathfrak{p}}\in\{\omega,\chi,\mathrm{ch}\}$ is different because for those parameters we have ${\mathfrak{p}}(k,n)\leq n+f(k,n)$ with either $f(k,n)=c(k)$ (i.e., the error term depends on $k$ but is independent of $n$) or $f(k,n)=c(k)\cdot h(n)$ where $h(n)$ is sublinear. On the other hand, our $\sqrt{k}\,n$ upper bound very quickly becomes substantially better than the corresponding estimates in the first three parts of Theorem 38.
Theorem 42.
$\omega(k,n)\leq\chi(k,n)\leq\mathrm{ch}(k,n)<\sqrt{k}\,n+k$.
8 Concluding remarks and open problems
In this section we present natural problems and directions for future research as a continuation suggested by the current research.
Problem 43.
Determine ${\mathfrak{p}}^{\mathsf{L}}(k,N)$ for graph parameters $\mathfrak{p}$} other than ${\mathfrak{p}}(G)=\mathrm{Mad}(G)$.
Problem 44.
Study the product version of ${\mathfrak{p}}(k,n)$ and ${\mathfrak{p}}^{\mathsf{L}}(k,N)$.
In cases where ${\mathfrak{p}}(k,n)$ is verified by a decomposition whose $\mathfrak{p}$}-values form a balanced sequence of equal or nearly equal integers, Jensen’s inequality is applicable and provides a strong upper bound. This is the case with the function ${\mathfrak{p}}^{\mathsf{L}}(k,N)$ as well:
$\max_{(G_{1},\ldots,G_{k})\ \mathrm{decompose}\ K_{n}}\Bigl\{\,\prod_{i=1}^{k}\mathrm{Mad}(G_{i})\Bigr\}\leq\Bigl(\frac{M(k,n)}{k}\Bigr)^{k}\leq\Bigl(\frac{M^{\mathsf{L}}(k,\binom{n}{2})}{k}\Bigr)^{k},$
$\max_{e(G_{1})+\ldots+e(G_{k})=N}\,\Bigl\{\,\prod_{i=1}^{k}\mathrm{Mad}(G_{i})\Bigr\}\leq\Bigl(\frac{M^{\mathsf{L}}(k,N)}{k}\Bigr)^{k}<\bigl(2N/k\bigr)^{k/2}.$
Our several results above show that this method works in certain ranges of $k,n$ for $M(k,n)$. For instance, the following holds.
Theorem 45.
Let $k=\binom{n}{2}/\binom{p}{2}$, and assume that a Steiner system $S(2,p,n)$ exists. Then
$\max_{(G_{1},\ldots,G_{k})\ \mathrm{decompose}\ K_{n}}\Bigl\{\,\prod_{i=1}^{k}\mathrm{Mad}(G_{i})\Bigr\}=(p-1)^{k}=(p-1)^{\frac{n(n-1)}{p(p-1)}}.$
Proof.
On applying part $(i)$ of Theorem 25 we see that in the present case M(k,n)/k=p1M(k,n)/k=p-1. The existence of Steiner systems ensures that the product of $\mathrm{Mad}$-values can now attain the universal upper bound $\bigl(M(k,n)/k\bigr)^{k}$. ∎
In a similar way, constructions from multisets $\{a*K_{p}\cup b*K_{p+1}\}$ yield exact ansewrs. However, this kind of tight constructions requires that $k$ is relatively large with respect to $n$. For $k$ small and $n$ large, the upper bound obtained via the geometric-arithmetic inequality is probably far from being tight, therefore the cases of fixed $k$ deserve attention. In particular, it is natural to ask:
Problem 46.
Determine  $\max\{\mathrm{Mad}(G)\cdot\mathrm{Mad}(\overline{G})\}$  over the graphs $G$ of order $n$. In particular, is this quantity maximized by a pair $G=K_{x}$,  $\overline{G}=K_{n}-K_{x}$, similarly to $\max\{\mathrm{Mad}(G)+\mathrm{Mad}(\overline{G})\}$ ?
Concerning ${\limsup_{n\to\infty}\max_{|G|=n}}\,\{\mathrm{Mad}(G)\cdot\mathrm{Mad}(\overline{G})/n^{2}\}$, Theorem 18 implies the upper bound $25/64=0.390625$, and the construction in its proof yields $3/8=0.375$ as lower bound on $\liminf$. We can improve on the latter, and obtain an estimate which is close to the $25/64$ bound within relative error less than $1.5$ %.
Proposition 47.
${\liminf_{n\to\infty}\max_{|G|=n}}\,\{\mathrm{Mad}(G)\cdot\mathrm{Mad}(\overline{G})/n^{2}\}\geq 2/\!\sqrt{27}>0.3849$.
Proof.
Let $G\cong K_{x}$, where $x\approx n/\!\sqrt{3}$. Then $\mathrm{Mad}(G)=x-1$, and $\mathrm{Mad}(\overline{G})=\bigl(x\cdot(n-x)+(n-x)\cdot(n-1)\bigr)/n\approx(n-x)(n+x)/n=n\cdot\bigl(1-(x/n)^{2}\bigr)\approx 2n/3$. ∎
Problem 48.
Improve on the lower bounds given in Proposition 19 for $3\leq k\leq 7$, or prove that those estimates are optimal.
Problem 49.
Extend the formula derived in Theorem 23 from the range $k\geq\frac{1}{6}\,n(n-1)$ to $k\geq\frac{1}{12}\,n(n-1)$.
Part of a solution would require the design of partial Steiner systems $PS(2,4,n)$ whose leave graphs admit edge decompositions into triangles. An $S(2,4,n)$ system exists if and only if $n\equiv 1$ or 4 (mod 12). However, a $\{K_{3},K_{4}\}$-design requires that $a$ triangles and $b$ quadruples incident with any one vertex satisfy 2a+3b=n12a+3b=n-1, which puts considerable arithmetic limitations on the combination of numbers of $K_{3}$ and $K_{4}$ subgraphs occurring in a decomposition. Therefore $M(k,n)$ in this range of $k$ will usually involve non-complete subgraphs in an optimal edge decomposition of $K_{n}$. We note that for $n\equiv 1$ (mod 12) there exist both $S(2,3,n)$ and $S(2,4,n)$ systems, starting from $n=13$. Also, $K_{10}$ decomposes into $\{9*K_{3},\,3*K_{4}\}$, and $K_{12}$ decomposes into $\{4*K_{3},\,9*K_{4}\}$.
Problem 50.
Characterize the extremal graphs for $M(2,n)$ determined in Theorem 18.
Some results of Section 3.2 maybe useful in this direction also.
Problem 51.
For ${\mathfrak{p}}\in\{\omega,\chi,\mathrm{ch}\}$ improve the $\sqrt{k}\,n$ upper bound in Theorem 42, and determine a strong lower bound on ${\mathfrak{p}}(k,n)$ as $k=f(n)$ grows not very slowly.
In particular, concerning $\omega$, we gratefully thank a referee for advising us the following observation, which expresses equivalence between the Zarankiewicz problem and the determination of $\omega(k,n)$.
Proposition 52.
For all $1\leq k\leq\binom{n}{2}$, $\omega(k,n)$ is equal to the largest number of edges in a $C_{4}$-free bipartite graph with $k$ and $n$ vertices in its partite sets.
Proof.
Assume first that there exists a $C_{4}$-free bipartite graph $B\subset K_{k,n}$ with $m$ edges, with vertex classes $\{a_{1},\dots,a_{k}\}$ and $\{b_{1},\dots,b_{n}\}$. Then the neighborhoods $N(a_{1}),\dots,N(a_{k})$ form a system of $k$ sets, any two of which share at most one vertex. Hence this can be viewed as an edge-disjoint packing of $k$ complete subgraphs in the complete $K_{n}$ with vertex set $\{b_{1},\dots,b_{n}\}$. Of course, they can be extended to an edge decomposition $G_{1},\dots,G_{k}$ of $K_{n}$. Consequently, $\omega(k,n)\geq\sum_{i=1}^{k}\omega(G_{i})\geq\sum_{i=1}^{k}|N(a_{i})|=m$.
To derive the reverse inequality, assume that $G^{\prime}_{1},\dots,G^{\prime}_{k}$ is a decomposition of $K_{n}$ with $V(K_{n})=\{b_{1},\dots,b_{n}\}$, such that $m^{\prime}:=\sum_{i=1}^{k}\omega(G^{\prime}_{i})=\omega(k,n)$. Choose a largest complete subgraph in $G^{\prime}_{i}$ ($i=1,\dots,k$), say on the vertex subset $X_{i}\subset V(G^{\prime}_{i})$, $|X_{i}|=\omega(G^{\prime}_{i})$. Since the $G^{\prime}_{i}$ are edge-disjoint, any two of the $X_{i}$ have at most one vertex in common. Let $B^{\prime}\subset K_{k,n}$ be the bipartite graph whose vertex classes are $\{a_{1},\dots,a_{k}\}$ and $\{b_{1},\dots,b_{n}\}$, and each $a_{i}$ has neighborhood $N(a_{i})=X_{i}$. This $B^{\prime}$ has $e(B^{\prime})=\sum_{i=1}^{k}|X_{i}|=\sum_{i=1}^{k}\omega(G_{i})=m^{\prime}=\omega(k,n)$, and it is $C_{4}$-free because each edge of $K_{n}$ belongs to only one $G_{i}$. ∎
As we noted earlier, in the quadratic range of $k$ one can improve $\sqrt{k}\,n$ to $(1-c)\sqrt{k}\,n$ with a suitably chosen constant $c>0$. But the gap between ${\mathfrak{p}}(k,n)$ and $M(k,n)$ may be considerably larger than that.
Based on Propositions 4 and 5, and the subsequent related observations under them, the following question is very natural to ask.
Problem 53.
Characterize the graphs for which $\mathrm{Mad}(G)=\mathrm{ed}(G)$ holds.
There are chances for a structural characterization because equality can be decided in polynomial time. Namely, it is immediate to compute $\mathrm{ed}(G)$ in linear time, and although definitely not so easy, also $\mathrm{Mad}(G)$ can be determined by a polynomial-time algorithm as designed by Goldberg [15]. On the other hand, a characterization in terms of forbidden induced subgraphs does not exist. This follows e.g. from the short discussion on the stackexchange.com site [24]. Upon a question posed by Cyriac Antony there, Tuukka Korhonen (a.k.a. Laakeri) commented that the complete join $H^{\prime}=H+\overline{K}_{|H|}$ satisfies $\mathrm{Mad}(H^{\prime})=d(H^{\prime})$ for every graph $H$.
Acknowledgement.
We thank the referees for their very constructive remarks and suggestions. Research of the second author was supported in part by the ERC Advanced Grant “ERMiD”.
Appendix: Alternative approach to Theorems 20 and 21
In this Appendix we provide an argument of algorithmic nature as an alternative proof of Theorem 21. It demonstrates that every decomposition of $K_{n}$ can be transformed step by step to a unique well-defined representative list of graphs. During this process the sum $\mathrm{Mad}(G_{1})+\ldots+\mathrm{Mad}(G_{k})$ never decreases, and finally yields the claimed upper bound on $M(k,n)$. The method is equally fine for a proof of Theorem 20, hence determining $M^{\mathsf{L}}(k,N)$ for arbitrary integers $N\geq k\geq 2$.
For convenience let us recall that the following will be proved:
Theorem 21. Let $n\geq 3$ and $2\leq k\leq\binom{n}{2}$ be given, and choose the integer $p$ for which $k\binom{p}{2}\leq\binom{n}{2}<k\binom{p+1}{2}$ holds. Let $q,r$ be the nonnegative uniquely determined integers such that $\binom{n}{2}=k\binom{p}{2}+qp+r$, where $0\leq q<k$ and $0\leq r<p$. Then
[…]
Moreover, equality holds whenever $K_{n}$ admits an edge decomposition into
  • $q$ copies of $K_{p+1}$ and kqk-q copies of $K_{p}$, if $r=0$; or
  • $q$ copies of $K_{p+1}$, kq1k-q-1 copies of $K_{p}$, and one copy of any one graph $G\in\mathcal{F}_{p,\,r}$, if $0<r<p$.
Proof.
Consider any edge decomposition $(G_{1},\dots,G_{k})$ of $K_{n}$. We forget that the $G_{i}$ originate from a decomposition; we only view them as a list of individual graphs, and manipulate with the graphs themselves. In this way, if we only assume that the starting collection of graphs is just a list $\mathsf{L}=\{G_{1},\dots,G_{k}\}$ with $e(G_{1})+\ldots+e(G_{k})=N$, then the process also leads to the determination of $M^{\mathsf{L}}(k,N)$ with the formula stated in Theorem 20. However, in the sequel we restrict attention to $N=\binom{n}{2}$.
1. We replace each $G_{i}$ with $H_{i}$, where $H_{i}$ is the extremal representative of the graphs with $e(G_{i})$ edges. Due to Theorem 8, we have $\mathrm{Mad}(H_{i})\geq\mathrm{Mad}(G_{i})$.
2. We qualify the graphs $H_{i}$ into three categories, where the involved integer $p$ is uniquely determined by the number $e(H_{i})$ of edges, and may vary for different subscripts among the $k$ graphs under consideration.
Type A:
$e(H_{i})=\binom{p_{i}}{2}$.
Type B:
$\binom{p_{i}}{2}<e(H_{i})<\binom{p_{i}+1}{2}$ and $\mathrm{Mad}(H_{i})=2e(H_{i})/(p_{i}+1)$.
Type C:
$\binom{p_{i}}{2}<e(H_{i})<\binom{p_{i}+1}{2}$ and $\mathrm{Mad}(H_{i})=p_{i}-1$.
At any stage of the procedure, we denote by $\mathcal{A}$, $\mathcal{B}$, $\mathcal{C}$ the actual collection (multiset) of graphs $H_{i}$ of type A, B, C, respectively. If $e(H_{i})=\frac{1}{2}(p_{i}+1)(p_{i}-1)=\binom{p_{i}}{2}+\frac{p_{i}-1}{2}$, then $H_{i}$ is qualified for both $\mathcal{B}$ and $\mathcal{C}$. In this case we assume $H_{i}\in\mathcal{B}$.
Beside these, we maintain an integer $s$ for the number of so-called “spare edges”, which is initialized as $s=0$.
3. The following steps are performed, keeping $s+\sum_{i=1}^{k}e(H_{i})$ unchanged throughout, hence maintaining the sum equal to $\binom{n}{2}$.
  1. 3.a
    First we eliminate the members of $\mathcal{C}$. If a $H_{i}\in\mathcal{C}$ occurs, say $H_{i}\cong G_{p_{i},\,r_{i}}$, we re-define $H_{i}:=K_{p_{i}}$ and $s:=s+r$. This keeps $\mathrm{Mad}(H_{i})$ unchanged, and increases $s$. In this way all $H_{i}\in\mathcal{C}$ are modified to type-A graphs at the beginning, and later on if a type-C graph arises, it is also eliminated immediately.
  2. 3.b
    Next we eliminate the members of $\mathcal{B}$, possibly with one exception. If $\mathcal{B}$ contains $H_{i},H_{j}$ such that $H_{i}\cong G_{p_{i},\,r_{i}}$ and $H_{j}\cong G_{p_{j},\,r_{j}}$, where either $p_{i}>p_{j}$, or $p_{i}=p_{j}$ and $r_{i}\leq r_{j}$, then transfer $r^{\prime}:=\min\{r_{i},p_{j}-r_{j}\}$ edges from $H_{i}$ to $H_{j}$. We do the same if $|\mathcal{B}|=1$, say $\mathcal{B}=\{H_{j}\}$ where $H_{j}=G_{p_{j},\,r_{j}}$, and $\mathcal{A}$ has a member $H_{i}\cong K_{p_{i}+1}=G_{p_{i},p_{i}}$ for some $p_{i}>p_{j}$.
    Note that either $r^{\prime}=p_{j}-r_{j}$ or $r^{\prime}=p_{j}-r_{j}-1$ holds, because the assumptions $H_{i},H_{j}\in\mathcal{B}$ and $p_{i}\geq p_{j}$ imply $p_{j}-r_{j}\leq(p_{j}+1)/2\leq(p_{i}+1)/2\leq r_{i}+1$. Hence, the following outcomes are possible after such a step.
    • If $r^{\prime}=r_{i}=p_{j}-r_{j}$, then $H_{i}\cong K_{p_{i}}$ and $H_{j}\cong K_{p_{j}+1}$, and both $H_{i}$ and $H_{j}$ have been moved from $\mathcal{B}$ to $\mathcal{A}$.
    • If $r^{\prime}=r_{i}=p_{j}-r_{j}-1$, then $H_{i}\cong K_{p_{i}}$ (moved from $\mathcal{B}$ to $\mathcal{A}$), and $H_{j}\cong G_{p_{j},p_{j}-1}$ (hence remains a member of $\mathcal{B}$).
    • If $r^{\prime}=p_{j}-r_{j}<r_{i}$, then $H_{j}\cong K_{p_{j}+1}$ (moved from $\mathcal{B}$ to $\mathcal{A}$), and $H_{i}\cong G_{p_{i},r_{i}-r^{\prime}}$, which either remains in $\mathcal{B}$ or has been moved from $\mathcal{B}$ to $\mathcal{C}$. In the latter case it is then immediately transformed to $K_{p_{i}}$ by step 3.a, hence becomes a member of $\mathcal{A}$.
    In either of these three cases, the value of $\mathrm{Mad}(H_{j})$ has been increased by exactly $2r^{\prime}\!/(p_{j}+1)$. On the other hand, the decrease of $\mathrm{Mad}(H_{i})$ is at most $2r^{\prime}\!/(p_{i}+1)$. Indeed, removing the $r^{\prime}$ edges from $H_{i}$ one by one, until reaching the threshold of $r_{i}=(p_{i}-1)/2$ the omission of each edge yields a loss of $2/(p_{i}+1)$, and no decrease occurs afterwards. Thus, the overall increase of $\mathrm{Mad}(H_{i})+\mathrm{Mad}(H_{j})$ is not smaller than $\frac{2(p_{i}-p_{j})r^{\prime}}{(p_{i}+1)(p_{j}+1)}$.
    In case $p_{i}=p_{j}$ it is possible that $\mathrm{Mad}(H_{i})+\mathrm{Mad}(H_{j})$ remains unchanged. This happens if the modified $H_{j}$ is $K_{p_{i}+1}$ while $H_{i}$ still remains a member of $\mathcal{B}$ (what means $r_{i}+r_{j}-p_{i}\geq(p_{i}-1)/2$).
    Performing 3.a and 3.b as long as possible, we obtain $\mathcal{C}=\varnothing$ and either $\mathcal{B}=\varnothing$ or $|\mathcal{B}|=1$.
  3. 3.c
    The next goal is to make a balance in $\mathcal{A}\cup\mathcal{B}$. Recall that $G_{p^{\prime},p^{\prime}}\cong K_{p^{\prime}+1}$ holds for every $p^{\prime}\geq 1$. Assume that $H_{i}=G_{p_{i},p_{i}}$ is the largest member of $\mathcal{A}$, and $H_{j}=G_{p_{j},p_{j}}$ is its smallest member; moreover if $\mathcal{B}\neq\varnothing$, let $\mathcal{B}=\{G_{p_{k},\,r_{k}}\}$ (where $0<r_{k}<p_{k}$). If $\max\{p_{i},p_{j},p_{k}\}>\min\{p_{i},p_{j},p_{k}\}+1$ (disregarding $p_{k}$ if $\mathcal{B}=\varnothing$), select the two of $H_{i},H_{j},H_{k}$ whose $p$-value is largest and smallest, respectively, and perform the transformation 3.b on them. (If the step is applied on $H_{i},H_{j}\in\mathcal{A}$, then $s$ is updated exactly as $s:=s+p_{i}-p_{j}-1$.) Repeat this until all $p$-values in $\mathcal{A}\cup\mathcal{B}$ differ by at most 1. In the same way as in the analysis of 3.b, it is readily seen that the sum of $\mathrm{Mad}$-values has been increased (and also $s$).
  4. 3.d
    In the last phase the $s$ spare edges are eliminated. If $\mathcal{B}\neq\varnothing$ and $H_{k}\cong G_{p_{k},\,r_{k}}$, first $r^{\prime\prime}:=\min\{s,p_{k}-r_{k}\}$ edges are inserted in $H_{k}$. The process is finished if $r^{\prime\prime}=s$. Otherwise we update $s:=s-r^{\prime\prime}$; at that moment $\mathcal{B}=\mathcal{C}=\varnothing$, and the family $\mathcal{A}$ as a multiset is of the form $\{a*K_{p^{\prime}},b*K_{p^{\prime}+1}\}$ (meaning $a$ copies of $K_{p^{\prime}}$ and $b$ copies of $K_{p^{\prime}+1}$) for some $p^{\prime}\geq 2$, $a\geq 1$, $b\geq 0$. As long as $s\geq p^{\prime}$, pick a copy of $K_{p^{\prime}}$ from $\mathcal{A}$, enlarge it to $K_{p^{\prime}+1}$, and update $s:=s-p^{\prime}$. Once this has been done for all $H_{i}\cong K_{p^{\prime}}$, and still $s>0$ holds, update $p^{\prime}:=p^{\prime}+1$ and continue in the same way.
  5. 3.e
    If $s=0$, stop. If $s>0$, replace one copy of $K_{p^{\prime}}$ with the graph $G_{p^{\prime},s}$. Call the current family $H_{1},\dots,H_{k}$ the representative of the pair $(k,n)$.
    Due to the choice of $p,q,r$ with respect to $k$, at this point $\mathcal{A}\cup\mathcal{B}\cup\mathcal{C}$ is of the form either $\{a*K_{p},b*K_{p+1},1*G_{p,r}\}$ with $1\leq r<p$ and $a,b\geq 0$ ($a=0$ or $b=0$ is possible, a+b=k1a+b=k-1), or $\{a*K_{p},b*K_{p+1}\}$ (here $a+b=k$, possibly with $b=0$), where $a,b$ depend on $n$ and $k$, and are easily expressible in terms of $q$ and $r$, as stated in the theorem.
Observe that, starting from an arbitrary edge decomposition of $K_{n}$ into $k$ subgraphs $G_{1},\dots,G_{k}$, the procedure terminates with the representative of the pair $(k,n)$ as the unique final family $H_{1},\dots,H_{k}$ determined above. Thus,
$M(k,n)\leq\mathrm{Mad}(H_{1})+\ldots+\mathrm{Mad}(H_{k})\,.$
The numerical formula follows from the facts that $\mathrm{Mad}(K_{p})=p-1$ (applied also as $\mathrm{Mad}(K_{p+1})=p$) and that $\mathrm{Mad}(G_{p,r})$ is either $p$ or $(p^{2}-p+2r)/(p+1)$, depending on whether $r\leq(p-1)/2$ or bigger. Namely, the procedure ends with the combination of graphs as displayed in the following table:
   graphs $H_{i}$ total of $\mathrm{Mad}(H_{i})$ side condition
$q$ copies of $K_{p+1}$ qpqp
kq1k-q-1 copies of $K_{p}$ (kq1)(p1)(k-q-1)(p-1)
11 copy of $G_{p,r}$ p1p-1 $r\leq(p-1)/2$
11 copy of $G_{p,r}$ p2(pr)/(p+1)p-2(p-r)/(p+1) r>(p1)/2r>(p-1)/2
Thus, summing up the entries in the middle column according to the side condition depending on the value of $r$, the claimed upper bound follows.
The cases of equality, as stated in the last part of the theorem, can also be verified. Indeed, if an assumed edge decomposition of $K_{n}$ exists, then all computations above hold with equality, and the proved upper bound on $M(k,n)$ is attained. ∎